### v1

# Provable Hyperparameter Tuning for Structured Pfaffian Settings

Abstract
Data-driven algorithm design automatically adapts algorithms to specific application domains, achieving better performance. In the context of parameterized algorithms, this approach involves tuning the algorithm’s parameters using problem instances drawn from the problem distribution of the target application domain. This can be achieved by maximizing empirical utilities that measure the algorithms’ performance as a function of their parameters, using problem instances.
While empirical evidence supports the effectiveness of data-driven algorithm design, providing theoretical guarantees for several parameterized families remains challenging. This is due to the intricate behaviors of their corresponding utility functions, which typically admit piece-wise and discontinuity structures. In this work, we present refined frameworks for providing learning guarantees for parameterized data-driven algorithm design problems in both distributional and online learning settings.
For the distributional learning setting, we introduce the Pfaffian GJ framework, an extension of the classical GJ framework, that is capable of providing learning guarantees for function classes for which the computation involves Pfaffian functions. Unlike the GJ framework, which is limited to function classes with computation characterized by rational functions, our proposed framework can deal with function classes involving Pfaffian functions, which are much more general and widely applicable. We then show that for many parameterized algorithms of interest, their utility function possesses a refined piece-wise structure, which automatically translates to learning guarantees using our proposed framework.
For the online learning setting, we provide a new tool for verifying dispersion property of a sequence of loss functions, a sufficient condition that allows no-regret learning for sequences of piece-wise structured loss functions where the piece-wise structure involves Pfaffian transition boundaries. We use our framework to provide novel learning guarantees for many challenging data-driven design problems of interest, including data-driven linkage-based clustering, graph-based semi-supervised learning, and regularized logistic regression.
1 Introduction
Data-driven algorithm design [GR16, Bal20] is a modern approach that develops and analyzes algorithms based on the assumption that problem instances come from an underlying application domain. Unlike traditional worst-case or average-case analyses, this approach leverages observed problem instances to design algorithms that achieve high performance for specific problem domains. In many application domains [BNVW17, BDSV18, BKST22], algorithms are parameterized, meaning they are equipped with tunable hyperparameters which significantly influence their performance. We develop general techniques applicable to a large variety of parameterized algorithm families for the selection of domain-specific good algorithms by learning the hyperparameters from the problem instances coming from the domain.
Typically, the performance of an algorithm is evaluated by a specific utility function. In more concrete terms, for an algorithm parameterized by $\bm{a}\in\mathcal{A}$, consider the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$, where $u_{\bm{a}}(\bm{x})$ gauges the performance of the algorithm with parameters $\bm{a}$ when inputting a problem instance $\bm{x}$. In the data-driven algorithm design setting, we assume an unknown underlying distribution $\mathcal{D}$ over $\mathcal{X}$, representing the application domain on which the algorithm operates. Consequently, designing algorithms tailored to a specific domain corresponds to the optimal selection of parameters ${\bm{a}}$ for the given domain.
Applications of data-driven algorithm design span various domains, including low-rank approximation [IVY19, IWW21, LLL+23], sparse linear systems solvers [LGM+20], dimensionality reduction [ALN21], among others. Its empirical success [IVY19, ACC+11, IWW21] underscores the necessity for a theoretical understanding of this approach. Intensive efforts have been made towards theoretical understanding for data-driven algorithm design, including learning guarantees for numerical linear algebra methods [BIW22], tuning regularization parameters for regression problems [BKST22, BNS24], unsupervised and semi-supervised learning [BDW18, BDL20, BS21], application in integer and mixed-integer programming [BDSV18, BPSV21], to name but a few.
Prior theoretical work on data-driven algorithm design focuses on two main settings: distributional (also known as statistical/batch) learning [BDD+21, BIW22] and online learning [BDV18]. In the distributional learning setting, there is a learner trying to optimize hyperparameters for the algorithm within a specific domain, given access to problem instances from that domain. In this case, the question is about the sample complexity: How many problem instances are required to learn good parameters that guarantee the algorithm’s performance in that domain? In the online learning setting, there is a sequence of problem instances chosen by an adversary arriving over time. The goal now is to design a no-regret learning algorithm: adjusting the algorithm’s parameters on the fly so that the difference between the average utility and the utility corresponding to the best parameters in hindsight diminishes over time.
The main challenge in establishing learning guarantees for the utility function classes lies in the complex structure of the utility function. In other words, even a very small perturbation in $\bm{a}$ can lead to a drastic change in the performance of the algorithm, making the analysis of such classes of utility functions particularly challenging. In response to this challenge, prior works take an alternative approach by analyzing the dual utility functions $\mathcal{U}^{*}=\{u_{\bm{x}}:\mathcal{A}\rightarrow[0,H]\mid\bm{x}\in\mathcal{X}\}$, which often admits piece-wise structured behavior [BDD+21, BNVW17, BIW22].
Building upon this observation, in the distributional learning setting, Balcan et al. [BDD+21] propose a general approach that analyzes the learnability of the utility function class via the piece and boundary functions class induced by the piece-wise structure. Targeting a more restricted scenario, Bartlett et al. [BIW22] introduced a refinement for the GJ framework [GJ93], which establishes tighter guarantees for utility function classes of which the piece-wise structure involves only rational functions. In the online learning setting, Balcan et al. [BDV18] introduce the dispersion condition, which serves as a sufficient condition allowing no-regret learning for piece-wise Lipschitz functions. Essentially, the dispersion property emphasizes that if the discontinuity of utility function sequences does not densely concentrate in any small region of parameter space, no-regret learning is possible. Despite their broad applicability, these generalized frameworks exhibit inherent limitations. In the distributional learning setting, the piece-wise structure framework introduced by Balcan et al. [BDD+21] often results in sub-optimal guarantees, especially when the parameters are multi-dimensional. Moreover, the framework requires analyzing the piece and boundary function classes as an intermediate step, which is not trivial and sometimes leads to loose or vacuous bounds. The refined GJ framework instantiated by Bartlett et al. [BIW22] is limited to the cases where the computation of utility functions only involves rational functions of the parameters. For the online learning setting, the dispersion property is generally challenging to verify [BDV18, BDP20, BS21], and requires many extra assumptions on the discontinuity of utility functions sequence. Moreover, when the form of discontinuity goes beyond affinity and rational functions, no prior verifying technique can be applied.
Motivated by the limitations of prior research, part of this work aims to present refined theoretical frameworks for data-driven algorithm design when the utility function admits a specific structure. In the distributional learning setting, we introduce a powerful Pfaffian GJ framework that can establish learning guarantees for function classes whose discontinuity involves Pfaffian functions. Roughly speaking, Pfaffian is a very general class of functions that captures a wide range of functions of interest, including rational, exponentiation, and combinations of those, among others. Furthermore, we demonstrate that many data-driven algorithm design problems exhibit a specific refined Pfaffian piece-wise structure, which, when combined with the Pfaffian GJ framework, can establish learning guarantees for such problems. In the online learning setting, we introduce a novel tool to verify the dispersion property, where the discontinuity of utility function sequences involves Pfaffian functions, which go beyond affinity and rational functions.
Another aim of this work is to provide learning guarantees for several under-investigated data-driven algorithm design problems, where the piece-wise structure of the utility functions involves Pfaffian functions. The problems we consider have been investigated in simpler settings, including data-driven agglomerative hierarchical clustering [BNVW17, BDL20], data-driven semi-supervised learning [BS21], and data-driven regularized logistic regression [BNS24]. However, previous investigations have limitations: they either have missing results for natural extensions of the settings under study [BNVW17, BS21], require strong assumptions [BDL20], or solely consider distributional learning settings [BNS24]. Moreover, we note that the techniques used in prior work are insufficient and cannot be applied in our settings, which involve Pfaffian analysis.
By carefully analyzing the utility functions associated with these problems, we uncover their underlying Pfaffian structures and control the Pfaffian complexities, which allows us to leverage our proposed frameworks to establish learning guarantees. It is important to note that analyzing those specific problems poses a significant challenge: a loose estimation of the Pfaffian function complexity when combined with our proposed frameworks would still lead to loose or vacuous learning guarantees.
Contributions.
In this work, we provide a refined framework for theoretical analysis of data-driven algorithm design problems. We then investigate many under-investigated data-driven algorithm design problems, analyzing their underlying problem structure, and then leveraging our newly proposed frameworks to provide learning guarantee for those problems. Concretely, our contributions are :
  • We present the Pfaffian GJ framework (5, Theorem 4.2), a general approach for analyzing the pseudo-dimension of various function classes of interest. This framework draws inspiration from the refined version of the GJ framework introduced by Bartlett et al. (2022) [BIW22]. However, in contrast to the conventional GJ framework which is only capable of handling computation related to rational functions, the Pfaffian GJ framework can handle computations that involve Pfaffian functions—a much broader function class—significantly increasing its applicability. We note that our proposed Pfaffian GJ framework is of independent interest and can be applied to other problems beyond data-driven algorithm design.
  • For distributional learning (statistical/batch) data-driven algorithm design, we introduce a refined piece-wise structure (8) for the dual utility function class, which applies whenever the piece and boundary functions are Pfaffian. In contrast to the prior piece-wise structure proposed by Balcan et al. [BDD+21], our framework assumes the piece and boundary functions belong to the broad class of Pfaffian functions and yields refined learning guarantees for this broad class of utility functions. We then show how the refined piece-wise structure can be combined with the newly proposed Pfaffian GJ framework to provide learning guarantees (Theorem 5.2) for problems that satisfy this property.
  • For online learning data-driven algorithm design, we introduce a general approach (Theorem 7.1, Theorem 7.2) for verifying the dispersion property [BDV18]—a sufficient condition for online learning in data-driven algorithm design. Prior work [BDP20, BS21] provides techniques for verifying dispersion only when the piece boundaries are algebraic functions. We significantly expand the class of functions for which online learning guarantees may be obtained by establishing a novel tool which applies for Pfaffian boundary functions.
  • We derive novel learning guarantees for a variety of understudied data-driven algorithm design problems, including data-driven agglomerative clustering (Theorem 6.1), data-driven graph-based semi-supervised learning (Theorem 6.4), as well as recover the guarantee for data-driven regularized logistic regression in previous work (Theorem D.3). By carefully analyzing the underlying structures of the utility functions for these problems, we can convert them into the form of our proposed frameworks, which automatically yield learning guarantees.
2 Related works
Data-driven algorithm design.
Data-driven algorithm design [GR16, Bal20] is a modern approach for automatically configuring algorithms and making them adaptive to specific application domains. In contrast to conventional algorithm design and analysis, which predominantly focuses on worst-case or average-case scenarios, data-driven algorithm design posits the existence of an (unknown) underlying problem distribution that dictates the problem instance that algorithms encounter. The primary objective shifts to identifying optimal configurations for such algorithms, leveraging available problem instances at hand from the same application domain. Empirical works have consistently validated the effectiveness of data-driven algorithm approaches in various domains, including matrix low-rank approximation [IVY19, IWW21], matrix sketching [LLL+23], mixed-integer linear programming [CKFB24], among others. These findings underscore the application of the data-driven algorithm design approach in real-world applications.
Distributional learning guarantee for data-driven algorithm design.
Motivated by the empirical success of data-driven algorithm design, there is an emerged line of work that focuses on theoretically analyzing the underlying mechanism, mostly focusing on providing generalization guarantees. This includes learning guarantee for data-driven algorithm designs of low-rank approximation and sketching [BIW22], learning metrics for clustering [BDL20, BS21, BDW18], integer and mixed-integer linear programming [BPSV21, BPSV22a, BDSV18], hyperparameter tuning for regularized regression [BKST22, BNS24], decision tree learning [BS24], robust nearest-neighbors [BBSZ23] among others.
Remarkably, Balcan et al. [BS21] introduced a general framework for establishing learning guarantees for problems that admit a specific piece-wise structure. Despite its broad applicability, the framework exhibits limitations: (1) it requires an intermediate task of analyzing the dual of piece and boundary functions, which is not trivial, and (2) it fails to incorporate extra useful information about piece and boundary functions, sometimes leading to sub-optimal bounds. Building on this insight, Bartlett et al. [BIW22] instantiates a refinement of the classical GJ framework [GJ93], offering an improved learning guarantee for data-driven algorithm design where the piece-wise structure involves only rational functions. However, it is essential to note that their framework cannot be applied beyond rational structures.
Online learning guarantee for data-driven algorithm design.
Another line of work focuses on providing no-regret learning guarantees for data-driven algorithm design problems in online learning settings. This includes online learning guarantees for heuristic knapsack, SDP-rounding for integer quadratic programming [BNVW17], and data-driven linkage-based clustering [BDP20], among others.
Online learning for data-driven algorithm design is generally a challenging task due to the discontinuity and piece-wise structure of the utility functions encountered. Most works provide learning guarantees in this setting by verifying the dispersion of the utility function sequence, a sufficient condition proposed by Balcan et al. [BDV18] stating that the discontinuity of the sequence is not highly concentrated in any small region. However, verifying the dispersion property is generally challenging. Balcan et al. [BDP20] provides a tool for verifying the dispersion property, targeting cases where the discontinuity is described by the roots of random polynomials for one-dimensional hyperparameters or algebraic curves in the two-dimensional case. Balcan and Sharma [BS21] then generalize this, providing a tool for discontinuities described by algebraic varieties in higher-dimensional cases.
Algorithms with predictions.
Another modern approach for design algorithms is algorithms with predictions [MV22], in which prediction is integrated at certain stages of algorithms to enhance their performance. Typically, this approach assumes the existence of a machine learning model capable of receiving problem instances as input and generating predictions specific to each instance. The performance analysis of these algorithms is closely tied to the quality of predictions generated by the machine learning models. A higher quality of predictions generally correlates with improved algorithmic performance. Hence, the algorithm, now integrated with predictive models, is analyzed based on its inherent algorithmic performance and the quality of predictions. algorithms with predictions have proved their efficacy in various classical problems, including support estimation [EIN+21], page migration [IMTMR22], online matching, flows, and load balancing [LMRX20], among others [KBTV22, LV21, WZ20].
While having many similarities and overlapping traits, there is a fundamental distinction between the two approaches. Data-driven algorithm design primarily seeks to optimize algorithmic parameters directly for specific application domains. On the other hand, algorithms with predictions. Besides, algorithms with predictions also have to decide what qualities or quantities of the problem instance to predict, and how to predict, which heavily affects the performance and properties of algorithms. It is worth noting that these two directions can complement each other and be integrated into the same system, as explored in numerous prior studies.
3 Preliminaries
We adhere closely to the problem setting and notation introduced by Balcan et al. [BDD+21]. Our focus is on parameterized algorithms, where the parameter belongs to a set $\mathcal{A}\subseteq\mathbb{R}^{d}$. Let $\mathcal{X}$ represent the set of input problem instances. The performance of the algorithm corresponding to parameter $\bm{a}\in\mathcal{A}$ on input $\bm{x}\in\mathcal{X}$ is given by the utility function $u(\bm{x},\bm{a})$ where $u:\mathcal{X}\times\mathcal{A}\rightarrow[0,H]$. The utility function class of that parameter is then defined as $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid a\in\mathcal{A}\}$. We then define the dual utility function class $\mathcal{U}^{*}=\{u^{*}_{\bm{x}}:\mathcal{A}\rightarrow[0,H]\mid\bm{x}\in\mathcal{X}\}$, where $u^{*}_{\bm{x}}(\bm{a}):=u(\bm{x},\bm{a})$, which consists of utility functions obtained by varying the parameter $\bm{a}$ for fixed problem instances from $\mathcal{X}$. The dual-class $\mathcal{U}^{*}$ plays an important role in our analysis, which we will discuss later.
Distributional learning.
In contrast to traditional worst-case or average-case algorithm analysis, we assume the existence of an underlying problem distribution $\mathcal{D}$ over $\mathcal{X}$, which encapsulates information about the relevant application domain. Under such an assumption, our goal is to determine how many problem instances are sufficient to learn the near-optimal parameters of the algorithm for such application-specific problem distribution. To this end, it suffices to bound the pseudo-dimension [Pol12] of the corresponding utility function class.
Definition 1 (Pseudo-dimension, [Pol12]).
Consider a real-valued function class $\mathcal{F}$, of which each function takes input in $\mathcal{X}$. Given a set of inputs $S=(\bm{x}_{1},\dots,\bm{x}_{N})$, we say that $S$ is shattered by $\mathcal{F}$ if there exists a set of real-valued threshold $r_{1},\dots,r_{N}\in\mathbb{R}$ such that |$\left|\{({\text{sign}}(f(\bm{x}_{1})-r_{1}),\dots,{\text{sign}}(f(\bm{x}_{N})-r_{N}))\mid f\in\mathcal{F}\}\right|=2^{N}$. The pseudo-dimension of $\mathcal{F}$, denoted as $\text{Pdim}(\mathcal{F})$, is the maximum size $N$ of a input set that $\mathcal{F}$ can shatter.
If the function class $\mathcal{F}$ is binary-valued, this corresponds to the well-known VC-dimension [VC74]. It is wide-known in traditional learning theory that a bound on the pseudo-dimension implies a bound on the generalization error (See Appendix A.1 for further background), which is the difference between the performance of the algorithm on the training problem instances and the expected performance over the application domain from which the instances are drawn.
Theorem 3.1 ([AB09]).
Given a real-valued function class $\mathcal{F}$ whose range is $[0,H]$, and assume that $\text{Pdim}(\mathcal{F})$ is finite. Then, given any $\delta\in(0,1)$, and any distribution $\mathcal{D}$ over the input space $\mathcal{X}$, with probability at least $1-\delta$ over the drawn of $S\sim\mathcal{D}^{n}$, we have
$\left|\frac{1}{n}\sum_{i=1}^{N}f(\bm{x}_{i})-\mathbb{E}_{\bm{x}\sim\mathcal{D}}[f(\bm{x})]\right|\leq O\left(H\sqrt{\frac{1}{N}\left(\text{Pdim}(\mathcal{F})+\ln\frac{1}{\delta}\right)}\right).$
Online learning.
In the online learning setting, there is a sequence of utility functions corresponds to a sequence of problem instances $(\bm{x}_{1},\dots,\bm{x}_{T})$, coming over $T$ rounds. The task is to design a sequence of parameters $(\bm{a}_{1},\dots,\bm{a}_{T})$ for the algorithm that minimize the expected regret regret
$\text{Regret}_{T}=\max_{\bm{a}\in\mathcal{A}}\sum_{t=1}^{T}u(\bm{x}_{t},\bm{a})-\sum_{t=1}^{T}u(\bm{x}_{t},\bm{a}_{t}).$
Our goal is to design a sequence of parameters $\bm{a}$ that achieve sub-linear regret.
4 Pfaffian GJ framework for data-driven algorithm design
In a classical work, Goldberg and Jerrum [GJ93] introduced a comprehensive framework for bounding the VC-dimension (or pseudo-dimension) of parameterized function classes exhibiting a specific property. They proposed that if any function within a given class can be computed via a specific type of computation, named a GJ algorithm, consisting of fundamental operators such as addition, subtraction, multiplication, division, and conditional statements, then the pseudo-dimension of such a function class can be effectively upper bounded. The bound depends on the running time of the algorithm, offering a convenient approach to reduce the task of bounding the complexity of a function class into the more manageable task of counting the number of operators.
However, a bound based on runtime can often be overly conservative. Recently, Bartlett et al. [BIW22] instantiated a refinement for the GJ framework. Noting that any intermediate values the GJ algorithm computes are rational functions of parameters, Bartlett et al. proposed more refined complexity measures of the GJ framework, namely the predicate complexity and the degree of the GJ algorithm. Informally, the predicate complexity and the degree are the number of distinct rational functions in conditional statements and the highest order of those rational functions, respectively. Remarkably, based on the refined complexity measures, Bartlett et al. showed a refined bound, demonstrating its efficacy in various cases, including applications on data-driven algorithm design for numerical linear algebra.
It is worth noting that the GJ algorithm has limitations as it can only accommodate scenarios where intermediate values are rational functions. Specifically, it does not capture more prevalent classes of functions, such as the exponential function. Building upon the insights gained from the refined GJ framework, we introduce an extended framework called the Pfaffian GJ Framework. Our framework can be used to bound the pseudo-dimension of function classes that can be computed not only by fundamental operators and conditional statements but also by a broad class of functions called Pfaffian functions, which includes exponential and logarithmic functions. Technically, our result is a refinement of the analytical approach introduced by [Kho91, KM97, MW97] which is directly applicable to data-driven algorithm design. An important part of our contribution is a careful instantiation of our main result for several important algorithmic problems, as a naive application could result in significantly looser bounds on the sample complexity.
4.1 Pfaffian functions
We present the foundational concepts of Pfaffian chains, Pfaffian functions, and their associated complexity measures. Introduced by Khovanskii [Kho91], Pfaffian function analysis is a tool for analyzing the properties of solution sets of Pfaffian equations. We note that these techniques have been previously used to derive an upper bound on the VC-dimension of sigmoidal neural networks [KM97].
We first introduce the notion of a Pfaffian chain. Intuitively, a Pfaffian chain consists of an ordered sequence of functions, in which the derivative of each function can be represented as a polynomial of the variables and previous functions in the sequence.
Definition 2 (Pfaffian Chain, [Kho91]).
A finite sequence of continuously differentiable functions $\eta_{1},\dots,\eta_{q}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and variables $\bm{a}=(a_{1},\dots,a_{d})\in\mathbb{R}^{d}$ form a Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$ if there are real polynomials $P_{i,j}(\bm{a},\eta_{1},\dots,\eta_{j})$ in $a_{1},\dots,a_{d},\eta_{1},\dots,\eta_{j}$, for for all $i\in[d]$ and $j\in[q]$, such that
$\frac{\partial\eta_{j}}{\partial a_{i}}=P_{i,j}(\bm{a},\eta_{1},\dots,\eta_{j}).$
We now define two complexity notations for Pfaffian chains, termed the length and Pfaffian degree, that dictate the complexity of a Pfaffian chain. The length of a Pfaffian chain is the number of functions that appear on that chain, while the Pfaffian degree of a chain is the maximum degree of polynomials that can be used to express the partial derivative of functions on that chain. Formal definitions of Pfaffian chain length and Pfaffian degree are mentioned in Definition 3.
Definition 3 (Complexity of Pfaffian chain).
Given a Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$, as defined in Definition 2, we say that the length of $\mathcal{C}$ is $q$, and Pfaffian degree of $\mathcal{C}$ is $\max_{i,j}\deg(P_{i,j})$.
Given a Pfaffian chain, one can define the Pfaffian function, which is simply a polynomial of variables and functions on that chain.
Definition 4 (Pfaffian functions, [Kho91]).
Given a Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$, as defined in Definition 2, a Pfaffian function over the chain $\mathcal{C}$ is a function of the form $g(\bm{a})=Q(\bm{a},\eta_{1},\dots,\eta_{q})$, where $Q$ is a polynomial in variables $\bm{a}$ and functions $\eta_{1},\dots,\eta_{q}$ in the chain $\mathcal{C}$.
The concepts of the Pfaffian chain, functions, and complexities may be a bit abstract to unfamiliar readers. To help readers better grasp the concepts of Pfaffian chains and Pfaffian functions, here are some simple examples.
Example 1.
Consider the chain $\mathcal{C}(a,e^{a})$ consisting of the variable $a$ and the function $e^{a}$, where $a\in\mathbb{R}$. Then $\mathcal{C}$ is a Pfaffian chain since $\frac{d}{da}e^{a}=e^{a}$, which is a polynomial of degree 1 in $e^{a}$. Hence, the chain $\mathcal{C}$ has length $q=1$ and Pfaffian degree $M=1$. Now, consider the function $f(a)=(e^{a})^{2}+a^{3}$. We observe that $f(a)$ is a polynomial in $a$ and $e^{a}$. Therefore, $f(a)$ is a Pfaffian function over the chain $\mathcal{C}$.
Example 2.
The following example is useful when analyzing the learnability of clustering algorithms. Let $\beta=(\beta_{1},\dots,\beta_{k})\in\mathbb{R}^{k}$ and $\alpha\in\mathbb{R}$ be variables, and let $d(\beta)=\sum_{i=1}^{k}d_{i}\beta_{i}>0$, where $d_{i}$ ($i\in[k]$) are some fixed real coefficients. Consider the functions $f(\alpha,\beta):=\frac{1}{d(\beta)}$, $g(\alpha,\beta):=\ln d(\beta)$, and $h(\alpha,\beta):=d(\beta)^{\alpha}$. Then $f$, $g$, and $h$ are Pfaffian functions from the chain $\mathcal{C}(\alpha,\beta,f,g,h)$ of length $3$ and Pfaffian degree $2$, since
$\frac{\partial f}{\partial\alpha}=0,$ $\frac{\partial f}{\partial\beta_{i}}=-d_{i}\cdot f^{2},$
$\frac{\partial g}{\partial\alpha}=0,$ $\frac{\partial g}{\partial\beta_{i}}=d_{i}\cdot f,$
$\frac{\partial h}{\partial\alpha}=g\cdot h,$ $\frac{\partial h}{\partial\beta_{i}}=d_{i}\cdot f\cdot h.$
4.2 Pfaffian GJ Algorithm
We now present a formal definition of the Pfaffian GJ algorithm, which shares similarities with the GJ algorithm but extends its capabilities to compute Pfaffian functions as intermediate values, in addition to basic operators and conditional statements. This improvement makes the Pfaffian GJ algorithm significantly more versatile compared to the classical GJ framework.
Definition 5 (Pfaffian GJ algorithm).
A Pfaffian GJ algorithm $\Gamma$ operates on real-valued inputs $\bm{a}\in\mathcal{A}\subseteq\mathbb{R}^{d}$, and can perform three types of operations:
[…]
  • Pfaffian operators of the form $v^{\prime\prime}=\eta(v)$, where $\eta:\mathbb{R}\rightarrow\mathbb{R}$ is a Pfaffian function.
  • Conditional statements of the form “if $\text{if }v\geq 0\dots\text{ else }\dots$”.
Here $v$ and $v^{\prime}$ are either inputs or (intermediate) values previously computed by the algorithm.
Definition 6 (Pfaffian chain associated with Pfaffian GJ algorithm).
Given a Pfaffian GJ algorithm $\Gamma$ operating on real-valued inputs $\bm{a}\in\mathcal{A}\subseteq\mathbb{R}^{d}$, we say that a Pfaffian chain $\mathcal{C}$ is associated with $\Gamma$ if all the intermediate values computed by $\Gamma$ is a Pfaffian function from the chain $\mathcal{C}$.
The main difference between the classical GJ algorithm [GJ93, BIW22] and the Pfaffian GJ algorithm is that the latter allows Pfaffian operators in its computation. By leveraging the fundamental properties of Pfaffian chains and functions, we can easily show that all intermediate functions computed by a specific Pfaffian GJ algorithm come from the same Pfaffian chain (see Appendix B.1 for details). This remarkable property enables us to control the complexity of the Pfaffian GJ algorithm by controlling the complexity of the corresponding Pfaffian chain. We formalize this claim in the following lemma.
Lemma 4.1.
For any Pfaffian GJ algorithm $\Gamma$ involving a finite number of operations, there is a Pfaffian chain $\mathcal{C}$ of finite length associated with $\Gamma$.
Proof Sketch. The Pfaffian chain $\mathcal{C}$ can be constructed recursively as follows. Initially, we create a Pfaffian chain of variables $\bm{a}$ with length 0. Using the basic properties of Pfaffian functions discussed in Section B.1, each time $\Gamma$ computes a new value $v$, one of the following cases arises: (1) $v$ is a Pfaffian function on the current chain $\mathcal{C}$, or (2) we can extend the chain $\mathcal{C}$ by adding new functions, increasing its length (but still finite), such that $v$ becomes a Pfaffian function on the modified chain $\mathcal{C}$. $\square$
Remark 1.
We note that each Pfaffian GJ algorithm $\Gamma$ can be associated with various Pfaffian chains, with different complexities. In any specific application, designing the corresponding Pfaffian chain $\mathcal{C}$ for $\Gamma$ with a small complexity is a crucial task that requires careful analysis.
We now present our main technical tool, which can be used to bound the pseudo-dimension of a function class by expressing the function computation a Pfaffian GJ algorithm and giving a bound on complexity of the associated Pfaffian chain. The detailed proof is presented in Appendix B.2.
Theorem 4.2.
Consider a real-valued function class $\mathcal{L}$ with domain $\mathcal{X}$, of which each algorithm $L_{\bm{a}}\in\mathcal{L}$ is parameterized by ${\bm{a}}\in\mathcal{A}\subseteq\mathbb{R}^{d}$. Suppose that for every $\bm{x}\in\mathcal{X}$ and $r\in\mathbb{R}$, there is a Pfaffian GJ algorithm $\Gamma_{\bm{x},r}$, with associated Pfaffian chain $\mathcal{C}_{\bm{x},r}$ of length at most $q$ and Pfaffian degree at most $M$, that given $L_{\bm{a}}\in\mathcal{L}$, check whether $L_{\bm{a}}(\bm{x})>r$. Moreover, assume that values computed at intermediate steps of $\Gamma_{\bm{x},r}$ are from the Pfaffian chain $\mathcal{C}_{\bm{x},r}$, each of degree at most $\Delta$; and the functions computed in the conditional statements are of at most $K$ Pfaffian functions. Then $\text{Pdim}(\mathcal{L})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta K+16d.$
Proof Sketch. The overall idea is that given $N$ input problem instances $\bm{x}_{1},\dots,\bm{x}_{N}$ and $N$ thresholds $r_{1},\dots,r_{N}$, we want to bound $\Pi_{\mathcal{L}}(N)$ the number of distinct sign patterns
$({\text{sign}}(L(\bm{x}_{N},\bm{a})-r_{1}>0),\dots,{\text{sign}}(L(\bm{x}_{N},{\bm{a}})-r_{N}>0))$
obtained by varying ${\bm{a}}\in\mathcal{A}$, where $L(\bm{x}_{i},{\bm{a}})=L_{\bm{a}}(\bm{x}_{i})$. Then we solve the inequality $2^{N}\leq\Pi_{\mathcal{L}}(N)$ to obtain an upper bound for $\text{Pdim}(\mathcal{L}).$
Based on the assumption, $\Pi_{\mathcal{L}}(N)$ can be upper bounded by the number of distinct sign patterns
$({\text{sign}}(\tau_{1}(\bm{x}_{1},{\bm{a}})-r_{1}),\dots,{\text{sign}}(\tau_{K}(\bm{x}_{1},{\bm{a}})-r_{1}),\dots,{\text{sign}}(\tau_{1}(\bm{x}_{N},{\bm{a}})-r_{N})),\dots,{\text{sign}}(\tau_{K}(\bm{x}_{N},{\bm{a}})-r_{N})).$
By using a result by Karpinski and Macintyre [KM97], we can turn the task above into the task of bounding the solution set connected components of a system of at most $d$ equations $\theta_{1}({\bm{a}})=0,\dots,\theta_{d}({\bm{a}})=0$, where $\theta_{i}:\mathcal{A}\rightarrow\mathbb{R}$ for $i=1,\dots,d$ is chosen from $\{r_{j}(\bm{x}_{k},\cdot)\}_{j\in[K],k\in[N]}$. Note that also from assumption, we can construct a Pfaffian chain of length $dq$ that contains all the Pfaffian functions $\theta_{i}$. Then, applying a result by Khovanskii [Kho91] yields the desired result. $\square$
Remark 2.
For the case $q=0$, meaning that the functions computed in the conditional statements are merely rational functions, Theorem 4.2 gives an upper bound of $\mathcal{O}(d\log(\Delta K))$, which matches the rate of GJ algorithm by Bartlett et al. [BIW22].
5 Pfaffian piece-wise structure for data-driven distributional learning
In this section, we propose the Pfaffian piece-wise structure for function classes, a refinement of the piece-wise decomposable structure introduced by Balcan et al. [BDD+21]. Compared to their piece-wise decomposable structure, our proposed Pfaffian piece-wise structure incorporates additional information about the Pfaffian structures of piece and boundary functions, as well as the maximum number of forms that the piece functions can take. We argue that the additional information can be derived as a by-product in many data-driven algorithm design problems, but being left out in the framework by Balcan et al. [BDD+21]. We then show that if the dual utility function class of a parameterized algorithm admits the Pfaffian piece-wise structure, we can establish an improved learning guarantee, compared to the framework by Balcan et al. [BDD+21], for the algorithm by leveraging our proposed Pfaffian GJ framework. Additionally, we propose a further refined argument for the case where all dual utility functions share the same boundary structures, which leads to further improved learning guarantee.
5.1 Prior generalization framework for piece-wise structured utility functions in data-driven algorithm design and its limitations
In this section, we discuss the utility of the Pfaffian GJ framework in providing learning guarantees for problems related to data-driven algorithm design. Many parameterized algorithms, such as combinatorial algorithms and integer/mixed-integer programming, exhibit volatile utility functions with respect to their parameters. In other words, even minor changes in parameters can lead to significant alterations in the behavior of the utility function. Analyzing such volatile utility function classes poses significant technical challenge.
Fortunately, many data-driven algorithm design problems still possess a certain degree of structure. Prior studies [BNVW17, BDSV18, BKST22, BNS24] have demonstrated that the dual utility functions associated with data-driven algorithm design problems often exhibit a piece-wise structure. In other words, the parameter space of the dual utility function can be partitioned into regions, each separated by distinct boundary functions. Within each region, the dual utility function corresponds to a piece function that exhibits well-behaved properties, such as being linear, rational, or Pfaffian in nature. Building upon this insight, work by Balcan et al. [BDD+21] proposed a formal definition of a piece-wise-structured dual utility function class and established a generalization guarantee applicable to any data-driven algorithm design problem conforming to such structures.
Formally, let us revisit the definition of the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$ for an parameterized algorithm, where $\mathcal{A}\subseteq\mathbb{R}^{d}$. This class represents functions that evaluate the performance of the algorithm, with $u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]$ denoting the utility function corresponding to parameter ${\bm{a}}$. For a given input problem instance $\bm{x}\in\mathcal{X}$, $u_{\bm{a}}(\bm{x})$ yields the performance evaluation of the algorithm on $\bm{x}$. Notably, for each input $\bm{x}\in\mathcal{X}$, we can define the dual utility function corresponding to $\bm{x}$ as $u_{\bm{x}}^{*}:\mathcal{A}\rightarrow[0,H]$, where $u^{*}_{\bm{x}}({\bm{a}}):=u_{\bm{a}}(\bm{x})$ measures the performance for a specific problem instance $\bm{x}$ as a function of the parameter ${\bm{a}}$. Consequently, we can also define the dual utility function class $\mathcal{U}^{*}=\{u^{*}_{\bm{x}}:\mathcal{A}\rightarrow\mathbb{R}\mid\bm{x}\in\mathcal{X}\}$. It was shown in prior works [BDD+21, BKST22, BPSV22b, BNS24] that in many data-driven algorithm design problems, every function in the dual function class $\mathcal{U}^{*}$ adheres to a specific piece-wise structure, which can be precisely defined as follows:
Definition 7 (Piece-wise decomposable, [BDD+21]).
A function class $\mathcal{H}\subseteq\mathbb{R}^{\mathcal{Y}}$ that maps a domain $\mathcal{Y}$ to $\mathbb{R}$ is $(\mathcal{F},\mathcal{G},k)$-piece-wise decomposable for a class $\mathcal{G}\subseteq\{0,1\}^{\mathcal{Y}}$ and a class $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{Y}}$ of piece functions if the following holds: for every $h\in\mathcal{H}$, there are $k$ boundary functions $g^{(1)},\dots,g^{(k)}\in\mathcal{G}$ and a piece function $f_{\mathbf{b}}\in\mathcal{F}$ for each bit vector $\{0,1\}^{k}$ such that for all $y\in\mathcal{Y}$, $h(y)=f_{\mathbf{b}_{y}}(y)$ where $\mathbf{b}_{y}=(g^{(1)}(y),\dots,g^{(k)}(y))\in\{0,1\}^{k}$.
A intuitive illustration of the piece-wise structure can be found in Figure 1. Roughly speaking, if a function class satisfies the piece-wise structure as defined in Definition 7, then for each function in such a class, the input domain is partitioned into multiple regions by $k$ boundary functions. Within each region, which corresponds to a $k$-element binary vector indicating its position relative to the $k$ boundary functions, the utility function is a well-behaved piece function. Based on this observation, Balcan et al. [BDD+21] showed that for any algorithm, if the dual utility function class satisfies the piece-wise structure as defined in Definition 7, then the pseudo-dimension of the utility function class is bounded.
Figure 1: An example of the original piece-wise structure (Definition 7) and our proposed Pfaffian piece-wise structure (Definition 8). In this example, fixed an input $x$, the parameter domain $\mathcal{A}$ is partitioned into six regions by three boundary functions $\mathbb{I}(g^{(1)}({\bm{a}})\geq 0)$, $\mathbb{I}(g^{(2)}({\bm{a}})\geq 0)$, and $\mathbb{I}(g^{(3)}({\bm{a}})\geq 0)$, where $\mathbb{I}$ denotes the indicator function. In each region, the utility function $u_{x}({\bm{a}})$ corresponds to a piece function $f_{\mathbf{b}}({\bm{a}})$ in which the binary vector $\mathbf{b}$ encodes the relative position of the region concerning the boundary functions. However, what is not captured by piece-wise structure is that: (1) the piece functions here can only take three analytical forms (each corresponding to a color), and (2) all the piece and boundary functions ($g^{(i)}$ and $f_{\mathbf{b}}$) are Pfaffian functions from the same Pfaffian chain.
Theorem 5.1 ([BDD+21]).
Consider the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$. Suppose that the dual function class $\mathcal{U}^{*}$ is $(\mathcal{F},\mathcal{G},k)$-piece-wise decomposable with boundary functions $\mathcal{G}\subseteq\{0,1\}^{\mathcal{A}}$ and piece functions $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{A}}$. Then the pseudo-dimension of $\mathcal{U}$ is bounded as follows
$\text{Pdim}(\mathcal{U})=\mathcal{O}((\text{Pdim}(\mathcal{F}^{*})+\text{VCdim}(\mathcal{G}^{*}))\log(\text{Pdim}(\mathcal{F}^{*})+\text{VCdim}(\mathcal{G}^{*}))+\text{VCdim}(\mathcal{G}^{*})\log k),$
where $\mathcal{F}^{*}$ and $\mathcal{G}^{*}$ is the dual class of $\mathcal{F}$ and $\mathcal{G}$, respectively.
Intuitively, Theorem 5.1 allows us to bound the pseudo-dimension of the utility function class $\mathcal{U}$, which is not well-behaved, by alternatively analyzing the dual boundary and piece function classes $\mathcal{F}^{*}$ and $\mathcal{G}^{*}$. However, Bartlett et al. [BIW22] demonstrate that for certain problems in which the piece-wise structure involves only rational functions, applying Theorem 5.1 can yield looser bounds compared to the approach based on the GJ framework [GJ93]. Furthermore, applying Theorem 5.1 requires the non-trivial task of analyzing the dual classes $\mathcal{F}^{*}$ and $\mathcal{G}^{*}$ of piece and boundary functions. This might cause trouble, such as leading to loose bounds (see [BDL20], Lemma 7) or vacuous bounds (see [BIW22], Appendix E.3). This situation arises when the piece and boundary functions involve Pfaffian functions, motivating the need to design a refined approach.
5.2 A refined piece-wise structure for data-driven algorithm design
In this section, we propose a more refined approach to derive learning guarantees for data-driven algorithm design problems where the utility functions exhibit a Pfaffian piece-wise structure. The key difference between our proposed frameworks and the framework by Balcan et al. [BDD+21] is that we consider the following additional factors: (1) Both piece and boundary functions are Pfaffian functions from the same Pfaffian chain, and (2) the maximum number of the distinct forms that the piece function can admit. Later, we will argue that by leveraging those extra structures, we can get a better pseudo-dimension upper bound by a logarithmic factor, compared to using the framework by [BDD+21]. The Pfaffian piece-wise structure is formalized as below.
Definition 8 (Pfaffian piece-wise structure).
A function class $\mathcal{H}\subseteq\mathbb{R}^{\mathcal{Y}}$ that maps domain $\mathcal{Y}\subseteq\mathbb{R}^{d}$ to $\mathbb{R}$ is said to be $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$ piece-wise structured if the following holds: for every $h\in\mathcal{H}$, there are at most $k_{\mathcal{G}}$ boundary functions of the forms $\mathbb{I}(g^{(1)}_{h}\geq 0),\dots,\mathbb{I}(g^{(k)}_{h}\geq 0)$, where $k\leq k_{\mathcal{G}}$, and a piece function $f_{h,\mathbf{b}}$ for each binary vector $\mathbf{b}\in\{0,1\}^{k}$ such that for all $y\in\mathcal{Y}$, $h(y)=f_{h,\mathbf{b}_{y}}(y)$ where $\mathbf{b}_{y}=(\mathbb{I}(g^{(1)}_{h}(y)\geq 0),\dots,\mathbb{I}(g^{(k)}_{h}(y)\geq 0))\in\{0,1\}^{k}$. Moreover, the piece functions $f_{h,\mathbf{b}}$ can take on of at most $k_{\mathcal{F}}$ forms, i.e. |$\left|\left\{f_{h,\bm{b}}\mid\bm{b}\in\{0,1\}^{k}\right\}\right|\leq k_{\mathcal{F}}$, and all the piece and boundary functions are Pfaffian functions of degree at most $\Delta$ over a Pfaffian chain $\mathcal{C}_{h}$ of length at most $q$ and Pfaffian degree at most $M$.
Figure 2: A demonstration of how the computation of a dual utility function satisfying Pfaffian piece-wise structure can be described by Pfaffian GJ algorithm. Given an input $\bm{x}\in\mathcal{X}$ and a threshold $r\in\mathbb{R}$, the function $u^{*}_{x}$ is piece-wise structured with boundary functions $g^{(i)}_{h}$ (for $i=1,\dots k$), and piece functions $f_{h,\mathbf{b}}$ ($\mathbf{b}\in\{0,1\}^{k}$). Note that, the piece functions $f_{h,\mathbf{b}}$ can take at most $k_{\mathcal{F}}$ forms and all the piece and boundary functions are Pfaffian functions from the chain $\mathcal{C}$.
An intuitive illustration of Pfaffian piece-wise structure and its comparison with the piece-wise structure by Balcan et al. [BDD+21] can be found in Figure 1. For a data-driven algorithm design problem with corresponding utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$ where $\mathcal{A}\subseteq\mathbb{R}^{d}$, we can see that if its dual utility function class $\mathcal{U}^{*}=\{u^{*}_{\bm{x}}:\mathcal{A}\rightarrow\mathbb{R}\mid\bm{x}\in\mathcal{X}\}$ admits the Pfaffian piece-wise structure as in Definition 8, then it can be computed using the Pfaffian GJ algorithm (see Figure 2 for a visualization). Therefore, we can use the established results for the Pfaffian GJ algorithm (Theorem 5) to derive learning guarantees for such problems. We formalize this claim in the following theorem.
Theorem 5.2.
Consider the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$ where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Suppose that the dual function class $\mathcal{U}^{*}=\{u^{*}_{\bm{x}}:\mathcal{A}\rightarrow[0,H]\mid\bm{x}\in\mathcal{X}\}$ is $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$-Pfaffian piece-wise structured, then the pseudo-dimension of $\mathcal{U}$ is bounded as follows
$\text{Pdim}(\mathcal{U})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta(k_{\mathcal{F}}+k_{\mathcal{G}})+16d.$
Remark 3.
The details of the differences between our refined Pfaffian piece-wise structure and the piece-wise structure by Balcan et al. [BDD+21] can be found in Appendix C.3. In short, compared to the framework by Balcan et al. [BDD+21], our framework offers: (1) an improved upper bound on the pseudo-dimension by a logarithmic factor (Theorem C.1), and (2) a more straightforward method for problems admitting a Pfaffian piece-wise structure, without invoking dual piece and boundary function classes, which are non-trivial to analyze. This might lead to loose bounds (see [BDL20], Lemma 7, or [BIW22], Appendix E.3).
5.3 Analyzing via approximation function class
In many applications, we might want to analyze the utility function class $\mathcal{U}$ indirectly by studying a surrogate utility function class $\mathcal{V}$ that is "sufficiently close" to $\mathcal{U}$. There are several reasons for this approach. For instance, if we do not know the closed form of $\mathcal{U}$, it becomes challenging to analyze its pseudo-dimension [BNS24]. In such cases, we may consider a surrogate function class $\mathcal{V}$ with a closed-form expression for convenience. Another scenario arises when the function class $\mathcal{U}$ is too complex, and analyzing a simpler surrogate function class $\mathcal{V}$ might yield better bounds [BSV20].
In this section, we consider a scenario where we approximate the parameterized utility function over a predefined partition $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ of the parameter space $\mathcal{A}$. Formally, a partition $\mathcal{P}$ of a set $\mathcal{A}$ is a collection $\{\mathcal{A}_{1},\dots,\mathcal{A}_{n}\}$ is a collection of non-empty subsets $\mathcal{A}_{i}$ of $\mathcal{A}$ that are pairwise disjoint and whose union is the entire set $\mathcal{A}$. This is a special case, as for every problem instance $\bm{x}$, the dual function $u_{\bm{x}}^{*}$ exhibits the same discontinuity structure, which can be leveraged to obtain a tighter bound. The following lemma highlights that if the dual function $v_{\bm{x}}^{*}$ exhibits the same discontinuity structure across all $\bm{x}$, and within each region, it is a Pfaffian function arising from a Pfaffian chain of bounded complexity, then the function class $\mathcal{U}$ admits a refined upper bound on its pseudo-dimension. The detailed proof is provided in Appendix C.2.
Lemma 5.3.
Consider a function class $\mathcal{V}=\{v_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$ where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Assume there is a partition $\mathcal{P}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{n}\}$ of the parameter space $\mathcal{A}$ such that for any problem instance $x\in\mathcal{X}$, the dual utility function $u^{*}_{x}$ is a Pfaffian function of degree at most $\Delta$ in region $\mathcal{A}_{i}$ from a Pfaffian chain $\mathcal{C}_{\mathcal{A}_{i}}$ of length at most $q$ and Pfaffian degree $M$. Then the pseudo-dimension of $\mathcal{V}$ is upper bounded as follows
$\text{Pdim}(\mathcal{V})=\mathcal{O}(q^{2}d^{2}+qd\log(\Delta+M)+qd\log d+\log n).$
Remark 4.
Essentially, Lemma 5.3 simplifies the analysis by restricting the complexity of the Pfaffian chain to individual regions $\mathcal{A}_{i}$, rather than across the entire partition $\mathcal{P}$. This often significantly reduces the length of the Pfaffian chain, which is the dominant term in the upper bound in most cases.
6 Applications of the Pfaffian piece-wise structure framework
In this section, we demonstrate how to leverage our proposed framework to establish new distributional learning guarantees, significantly expanding the scope of algorithmic families data-driven algorithm design problems.
6.1 Data-driven agglomerative hierarchical clustering
Agglomerative hierarchical clustering (AHC) [MC12] is a versatile, two-stage clustering approach widely employed across various domains. In the first stage, data is organized into a hierarchical clustering tree, determining the order in which data points are merged into clusters. Subsequently, in the second stage, the cluster tree is pruned according to a specific objective function, of which some common choices are $k$-mean  $k$-median, and $k$-center [Llo82, XW05], among others, to obtain the final clusters.
The first stage of AHC involves carefully designing linkage functions, which measure the similarity between clusters and determine the pair of clusters to be merged at each step. These linkage functions require a pairwise distance function $d$ between data points and calculate the distance between clusters based on the distances of their constituent points in a specific manner. Common linkage functions include single-linkage, complete-linkage, and average-linkage, with numerous variants interpolating between these simple functions employed in practical applications [ABV17, SMB+03, WNN+10]. It is important to note that if two linkage functions generate the same hierarchical cluster tree, they will yield the same final clusters, irrespective of the objective function used in the subsequent pruning stage.
Although linkage functions are a crucial component of AHC, they are generally chosen heuristically without any theoretical guarantees. Recently, Balcan et al. [BNVW17] proposed a data-driven, provable approach for designing linkage functions. Similar to other data-driven settings, their analysis operates under the assumption that there exists an unknown, application-specific distribution for clustering instances. They then provide learning guarantees for some simple families of linkage functions, parameterized by a single parameter, that interpolates among single, complete, and average linkage. However, they assume that the pairwise distance function $d$ is fixed, whereas in practice, multiple distance functions, each with distinct properties and benefits, are combined to achieve better performance [BBC+05]. Subsequent work by Balcan et al. [Bal20] proposes combining multiple pairwise distance functions by jointly learning their weights and the parameters of the linkage function. However, their analysis holds only under a strict assumption on the linkage function and distance functions. In particular, it requires that the merge function $m$ is 2-point-based, meaning that the distance $m_{\delta}(A,B)$ between two clusters $A,B$ only depends on the distance $\delta(a,b)$ between two points $a\in A$ and $b\in B$. Moreover, that pair of points must depend only on the ordering of pairwise distances, in the sense that for any two distances $\delta,\delta^{\prime}$, and two pair of points $(a,b),(a^{\prime},b^{\prime})\in A\times B$, we have $\delta(a,b)\leq\delta(a^{\prime},b^{\prime})$ iff $\delta^{\prime}(a,b)\leq\delta^{\prime}(a^{\prime},b^{\prime})$, then $m_{\delta}(A,B)=\delta(a,b)$ implies that $m_{\delta^{\prime}}(A,B)=\delta^{\prime}(a,b)$. That restriction effectively rules out many merge functions of interests and compatibility of distance functions learned.
In this section, we instantiate theoretical guarantees for novel data-driven AHC settings in which we near-optimally learn the parameter of the merge function family and the combination of multiple point-wise distance functions. Our analysis requires no restriction on the considered point-wise distances, extends to cluster families that are not 2-point-based, and applies to any objective functions used for tree-pruning.
6.1.1 Problem setting
Given a set of $n$ points $S\in\mathcal{X}^{n}$, where $\mathcal{X}$ denotes the data domain, and a distance function $\delta:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$, the goal is to partition $S$ into clusters such that the intra-cluster distance is minimized, and the inter-cluster distance is maximized. In the AHC approach, we first design a linkage function $m_{\delta}$ based on $\delta$, where $m_{\delta}(A,B)$ specifies the distance between two clusters $A$ and $B$. The cluster tree construction algorithm starts with $n$ singleton clusters, each containing a single data point, and successively merges the pair of clusters $A,B$ that minimizes the cluster-wise distance $m_{\delta}(A,B)$. This sequence of merges yields a hierarchical cluster tree, with the root corresponding to the entire set $S$, leaf nodes corresponding to individual points in $S$, and each internal node representing a subset of points in $S$ obtained by merging the point sets corresponding to its two child nodes. Subsequently, the cluster tree is pruned to obtain the final clusters via a dynamic programming procedure that optimizes a chosen objective function. Common objective functions include $k$-means, $k$-medians, and $k$-centers, among others. Importantly, given a fixed objective function, if two linkage functions generate the same cluster tree for a given dataset, they will yield the same final clusters.
As discussed previously, the point-wise distance function $\delta$ can be a convex combination of several distance functions chosen from $\bm{\delta}=\{\delta_{1},\dots,\delta_{L}\}$, i.e., $\delta=\delta_{\beta}=\sum_{i=1}^{L}\beta_{i}\delta_{i}$, where $\beta=(\beta_{1},\dots,\beta_{L})\in\Delta(L)$. Here $\Delta(L)$ denotes the $(L-1)$-dimensional simplex. The combined distance function $\delta_{\beta}$ is then used in the linkage function. In this work, we consider the following parameterized families of linkage functions:
$\mathcal{M}_{1}$ =$=\Bigg{\{}m^{1,\alpha}_{\delta}:(A,B)\mapsto\left(\min_{a\in A,b\in B}(\delta(a,b))^{\alpha}+\max_{a\in A,b\in B}(\delta(a,b))^{\alpha}\right)^{1/\alpha},\alpha\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\Bigg{\}},$
$\mathcal{M}_{2}$ =$=\Bigg{\{}m^{2,\alpha}_{\delta}:(A,B)\mapsto\left(\frac{1}{|A||B|}\sum_{a\in A,b\in B}(\delta(a,b))^{\alpha}\right)^{1/\alpha},\alpha\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\Bigg{\}},$
$\mathcal{M}_{3}$ =$=\Bigg{\{}m^{3,\alpha}_{\delta}:(A,B)\mapsto\left(\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}\right)^{1/\sum_{i}\alpha_{i}},\alpha_{i}\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\Bigg{\}}.$
The merge function family $\mathcal{M}_{1}$ interpolates between single and complete linkage. The versatile linkage $\mathcal{M}_{2}$ is another rich linkage family, which interpolates among single, complete, and average linkage [FG20]. The family $\mathcal{M}_{3}$ is another generalization of single, complete, and average linkages that incorporates multiple pairwise distance functions. In $\mathcal{M}_{3}$, we do not combine all the given distance functions $\bm{\delta}=\{\delta_{1},\dots,\delta_{L}\}$ into one but treat them separately. Precisely, if we set $\alpha_{i}=1$, and $\alpha_{j}=0$ for all $j\neq i$, we have average linkage with respect to the distance function $\delta_{i}$; if we set $\alpha_{i}=\infty$, and $\alpha_{j}=0$ for all $j\neq i$, we have complete linkage with respect to the distance function $\delta_{i}$; and if we set $\alpha_{i}=-\infty$, and $\alpha_{j}=0$ for $j\neq i$, we have the well-known single linkage with respect to $\delta_{i}$.
In the data-driven setting, we are given multiple problem instances $P_{1},\dots,P_{N}$, where each instance $P_{i}$ consists of a set of points $S$ and a set of distance functions $\bm{\delta}$. Assuming that there exists an underlying problem distribution that represents a specific application domain, we aim to determine how many problem instances are sufficient to learn the parameters of the linkage function families and the weights of the distance functions (for the families $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$). These questions are equivalent to analyzing the pseudo-dimension of the following classes of utility functions.
For $i\in{1,2}$, let $A^{\alpha,\beta}_{i}$ denote the algorithm that takes a problem instance $P=(S,\bm{\delta})$ as input and returns a cluster tree $A^{\alpha,\beta}_{i}(S,\bm{\delta})\in\mathcal{T}$, where $\mathcal{T}$ is the set of all possible cluster trees, by using the interpolated merge function $m^{1,\alpha}_{\delta_{\beta}}$. Then given an objective function, for example the $k$-means objective, the pruning function $\mu_{k}:\mathcal{T}\rightarrow\mathcal{S}_{k}$ takes as input a clustering tree, and returns a $k$-partition $\{\mathcal{P}_{1},\dots,\mathcal{P}_{k}\}$ of $S$ that minimizes such objective function. Then, given a target cluster $\mathcal{Y}=\{C_{1},\dots,C_{k}\}$, the performance of the algorithm $A_{i}^{\alpha,\beta}$ is given by the Hamming distance between the produced clusters $\mu_{k}(A_{i}^{\alpha,\beta}(S,\bm{\delta}))=\{\mathcal{P}_{1},\dots,\mathcal{P}_{k}\}$ and the target clusters $\mathcal{Y}=\{C_{1},\dots,C_{k}\}$
$\ell(\mu_{k}(A_{i}^{\alpha,\beta}(S,\bm{\delta})),\mathcal{Y})=\min_{\sigma\in\mathbb{S}_{k}}\frac{1}{|S|}\sum_{i=1}^{k}|C_{i}\setminus P_{\sigma_{i}}|.$
Here, the minimum is taken over the set of all permutations of $\{1,\dots,k\}$. We can clearly see that $\ell$ takes value in $\{0,\frac{1}{n},\dots,\frac{n-1}{n},1\}$. However, note that given an objective function, the cluster tree is equivalent to the produced clusters. Thus, the performance of the algorithm is completely determined by the cluster tree it produces, and for simplicity, we can express the performance of $A^{\alpha,\beta}_{i}$ as $u^{\alpha,\beta}_{i}:(S,\bm{\delta})\mapsto v(A^{\alpha,\beta}_{i}(S,\bm{\delta}))$, where $v$ is a function that maps a cluster tree to a value in $[0,1]$.
Similarly, we denote $A_{3}^{\alpha}$ as the cluster tree building algorithm that takes $P=(S,\bm{\delta})$ as the input and returns a cluster tree $A_{3}^{\alpha}(S,\bm{\delta})$ by using the linkage function $m_{d}^{3,\alpha}$. Again, we have that $u_{3}^{\alpha}:(S,\bm{\delta})\mapsto v(A_{3}^{\alpha}(S,\bm{\delta}))$ represents the performance of the algorithm. We consider the following function classes that measure the performance of the above algorithm families:
$\mathcal{H}_{1}=\{u^{\alpha,\beta}_{1}:(S,\bm{\delta})\mapsto u(A^{\alpha,\beta}_{1}(S,\bm{\delta}))\mid\alpha\in\mathbb{R}\cup\{-\infty,+\infty\},\beta\in\Delta([L])\},$
$\mathcal{H}_{2}=\{u^{\alpha,\beta}_{2}:(S,\bm{\delta})\mapsto u(A^{\alpha,\beta}_{2}(S,\bm{\delta}))\mid\alpha\in\mathbb{R}\cup\{-\infty,+\infty\},\beta\in\Delta([L])\},$
$\mathcal{H}_{3}=\{u^{\alpha}_{3}:(S,\bm{\delta})\mapsto u(A^{\alpha}_{3}(S,\bm{\delta}))\mid\alpha_{i}\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\}.$
In the next section, we analyze the pseudo-dimension of the function classes described above, which provides insights into the number of problem instances required to learn near-optimal AHC parameters for a specific application domain.
6.1.2 Generalization guarantees for data-driven hierarchical clustering
In this section, we will leverage our proposed framework (Theorem 5.2) to establish the upper bounds for the pseudo-dimension of $\mathcal{H}_{1}$ described above.
Theorem 6.1.
Let $\mathcal{H}_{1}$ be a class of functions
$\mathcal{H}_{1}=\{u^{\alpha,\beta}_{1}:(S,\bm{\delta})\mapsto u(A^{\alpha,\beta}_{1}(S,\bm{\delta}))\mid\alpha\in\mathbb{R}\cup\{-\infty,+\infty\},\beta\in\Delta([L])\}$
mapping clustering instances $(S,\bm{\delta}))$ to $[0,1]$ by using merge functions from class $\mathcal{M}_{1}$ and an arbitrary objective function. Then $\text{Pdim}(\mathcal{H}_{1})=\mathcal{O}(n^{4}L^{2})$.
Proof.
The high-level idea is to show that the utility function exhibits a piece-wise structure: its parameter space is partitioned by multiple boundary functions, and within each region, the cluster tree remains unchanged, implying that the utility function is constant. We then characterize the number and complexity of the boundary functions, which we show belong to a Pfaffian system. Subsequently, we can utilize our framework to obtain a bound on the pseudo-dimension of $\mathcal{H}_{1}$.
Fix a problem instance $(S,\bm{\delta})$. Suppose $A,B\subseteq S$ and $A^{\prime},B^{\prime}\subseteq S$ are two candidate clusters at some merge step of the algorithm. Then $A,B$ is preferred for merging over $A^{\prime},B^{\prime}$ iff
$\left(\min_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}+\max_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}\right)^{1/\alpha}\leq\left(\min_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}+\max_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}\right)^{1/\alpha},$
or equivalently,
$(\delta_{\beta}(a_{1},b_{1}))^{\alpha}+(\delta_{\beta}(a_{2},b_{2}))^{\alpha}\leq(\delta_{\beta}(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha}+(\delta_{\beta}(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha},$
where we have set $a_{1},b_{1}=\operatorname*{arg\,min}_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}$, $a_{2},b_{2}=\operatorname*{arg\,max}_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}$, $a^{\prime}_{1},b^{\prime}_{1}=\operatorname*{arg\,min}_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}$, and $a^{\prime}_{2},b^{\prime}_{2}=\operatorname*{arg\,max}_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}$. For distinct choices of the points $a_{1},b_{1},a_{2},b_{2},a^{\prime}_{1},b^{\prime}_{1},a^{\prime}_{2},b^{\prime}_{2}$, we get at most $n^{8}$ distinct boundary conditions across which the merge decision at any step of the algorithm may change.
We next show that the boundary functions constitute a Pfaffian system in $\alpha,\beta_{1},\dots,\beta_{L}$ and bound its complexity. For each pair of points $a,b\in S$, define $f_{a,b}(\alpha,\beta):=\frac{1}{\delta_{\beta}(a,b)}$, $g_{a,b}(\alpha,\beta):=\ln{\delta_{\beta}(a,b)}$ and $h_{a,b}(\alpha,\beta):={\delta_{\beta}(a,b)}^{\alpha}$. Recall $\delta_{\beta}(a,b)=\sum_{i=1}^{L}\beta_{i}\delta_{i}(a,b)$. Consider the chain $\mathcal{C}(\alpha,\beta,f_{a,b},g_{a,b},h_{a,b})$ of length $q=3n^{2}$, for $a,b\in S$. We will show that these functions form a Pfaffian chain of Pfaffian degree $M=2$. Indeed, we have for each $a,b\in S$,
$\frac{\partial f_{a,b}}{\partial\alpha}=0,$ $\frac{\partial f_{a,b}}{\partial\beta_{i}}=-\delta_{i}(a,b)f_{a,b}^{2},$
$\frac{\partial g_{a,b}}{\partial\alpha}=0,$ $\frac{\partial g_{a,b}}{\partial\beta_{i}}=\delta_{i}(a,b)f_{a,b},$
$\frac{\partial h_{a,b}}{\partial\alpha}=g_{a,b}h_{a,b},$ $\frac{\partial g_{a,b}}{\partial\beta_{i}}=\delta_{i}(a,b)f_{a,b}h_{a,b},$
which establishes the claim. The boundary conditions can clearly be all written in terms of the functions $\{h_{a,b}\}_{a,b\in S}$, meaning that the degree $\Delta=1$. Note that the piece functions are constant and can only take value in $\{0,\frac{1}{n},\dots,1\}$, meaning that $k_{\mathcal{F}}=n+1$. Then we conclude that $\mathcal{H}_{1}^{*}$ admits $(n+1,n^{8},3n^{2},2,1,L+1)$-Pfaffian piece-wise structure. Applying Theorem 5.2 now gives that $\text{Pdim}(\mathcal{H}_{1})=\mathcal{O}(n^{4}L^{2}+n^{2}L\log L+L\log(n^{8}+n+1))=\mathcal{O}(n^{4}L^{2})$. ∎
Similarly, we also have the upper-bound for the pseudo-dimension of $\mathcal{H}_{2}$, and $\mathcal{H}_{3}$.
Theorem 6.2.
Let $\mathcal{H}_{2}$ be a class of functions
$\mathcal{H}_{2}=\{u^{\alpha,\beta}_{2}\colon(S,\bm{\delta})\mapsto u(A^{\alpha,\beta}_{2}(S,\bm{\delta}))\mid\alpha\in\mathbb{R}\cup\{-\infty,+\infty\},\beta\in\Delta([L])\}$
mapping clustering instances $(S,\bm{\delta})$ to $[0,1]$ by using merge functions from class $\mathcal{M}_{2}$ and an arbitrary merge function. Then $\text{Pdim}(\mathcal{H}_{2})=\mathcal{O}(n^{4}L^{2})$.
Theorem 6.3.
Let $\mathcal{H}_{3}$ be a class of functions
$\mathcal{H}_{3}=\{u^{\alpha}_{3}:(S,\bm{\delta})\mapsto u(A^{\alpha}_{3}(S,\bm{\delta}))\mid\alpha_{i}\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\}$
mapping clustering instances $(S,\bm{\delta})$ to $[0,1]$ by using merge functions from class $\mathcal{M}_{3}$. Then $\text{Pdim}(\mathcal{H}_{3})=\mathcal{O}(n^{4}L^{2})$.
The detailed proofs of Theorem 6.2, 6.3 for the function classes $\mathcal{H}_{2}$, and $\mathcal{H}_{3}$ can be found in Appendix D.1. Although these function classes share the same pseudo-dimension upper-bound, their structures differ, necessitating separate analyses and leading to distinct Pfaffian piece-wise structures. To show that the dual function classes of $\mathcal{H}_{1}$, $\mathcal{H}_{2}$, and $\mathcal{H}_{3}$ admit Pfaffian piece-wise structure, we analyze the transition condition on the hyperparameters when the preference for merging two candidate clusters $A,B$ switches to merging a different pair of clusters $A^{\prime},B^{\prime}$ instead, at any merge step of the hierarchical clustering algorithm. The transition condition corresponds to an equality involving Pfaffian functions of the parameters $\alpha$ and $\beta$. All of such equations corresponding to each tuple $(A,B,A^{\prime},B^{\prime})\subset S^{4}$ will divide the parameter space into regions, in each of which the AHC algorithm produces the same clustering tree, leading to the same performance. After this step, we construct the Pfaffian chains for each function in function classes $\mathcal{H}_{1}$, $\mathcal{H}_{2}$, and $\mathcal{H}_{3}$, where the difference naturally lies in the form of functions in those function classes. We then carefully analyze the complexities of the Pfaffian chain of those Pfaffian functions to obtain the above bounds.
6.2 Data-driven graph-based semi-supervised learning
Semi-supervised learning [CSZ10] is a learning paradigm where labeled data is scarce due to expensive labeling processes. This paradigm leverages unlabeled data in addition to a small set of labeled samples for effective learning. A common semi-supervised learning approach is the graph-based method [ZG09, CSZ10], which captures relationships between labeled and unlabeled data using a graph. In this approach, nodes represent data points, and edges are constructed based on the similarity between data point pairs, measured by a given distance function. Optimizing a labeling function on this graph helps propagate labels from the labeled data to the unlabeled data.
A large body of research focuses on how to learn such labeling functions given the graph, including using $st$-mincuts [BC01], optimizing harmonic objective with soft mincuts [ZGL03], label propagation [ZG02], among others. However, under the assumption that the graph well describes the relationship amongst data, it is known that all algorithms for learning the labeling function above are equivalent [ZG09]. This also highlights the importance of the graph in graph-based semi-supervised learning.
Despite its significance, the graph is usually considered given or constructed using heuristic methods without theoretical guarantees [Zhu05, ZCP04]. Recently, Balcan and Sharma [BS21] proposed a novel data-driven approach for constructing the graph, by learning the parameters of the graph kernel underlying the graph construction, from the problem instances at hand. Each problem instance $P$ consists of sets of labeled $\mathcal{L}$ and unlabeled data $\mathcal{U}$ and a distance metric $d$. Assuming that all problem instances are drawn from an underlying, potentially unknown distribution, they provide guarantees for learning near-optimal graph kernel parameters for such a distribution.
Nonetheless, they consider only a single distance function, whereas in practical applications, combining multiple distance functions, each with its unique characteristics, can improve the graph quality and typically result in better outcomes compared to utilizing a single metric [BBC+05]. In this section, we consider a novel and more practical setting for data-driven graph-based semi-supervised learning, where we learn the parameters of the commonly-used Gaussian RBF kernel $w_{\sigma,\beta}(u,v)=\exp(-\delta(u,v)/\sigma^{2})$ and the weights $\beta$ of $\delta=\sum_{i=1}^{L}\beta_{i}\delta_{i}$ is a combination multiple distance functions for constructing the graph.
6.2.1 Problem settings
In the graph-based semi-supervised learning with Gaussian RBF kernel, we are given a few labeled samples $\mathcal{L}\subset\mathcal{X}\times\mathcal{Y}$, a large number of unlabeled points $\mathcal{U}\subset\mathcal{X}$, and a set of distance functions $\bm{\delta}={\delta_{1},\dots,\delta_{L}}$, where $\delta_{i}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}{\geq 0}$ for $i\in[L]$. Here, $\mathcal{X}$ denotes the data space, and $\mathcal{Y}={0,1}$ denotes the binary classification label space. To extrapolate labels from $\mathcal{L}$ to $\mathcal{U}$, a graph $G^{\sigma,\beta}$ is constructed with the node set $\mathcal{L}\cup\mathcal{U}$ and edges weighted by the Gaussian RBF graph kernel $w_{\sigma,\beta}(u,v)=\exp(-d(u,v)/\sigma^{2})$, where $\sigma$ is the bandwidth parameter, and $\delta=\sum_{i=1}^{L}\beta_{i}\delta_{i}$ is a convex combination of the given distance functions. After constructing the graph $G^{\sigma,\beta}$, a popular graph labeling algorithm called the harmonic method [ZGL03] is employed. It assigns soft labels by minimizing the following quadratic objective:
$\frac{1}{2}\sum_{u,v}w(u,v)(f(u)-f(v))^{2}=f^{T}(D-W)f,$
where $f\in[0,1]^{n}$, $n=|\mathcal{U}|+|\mathcal{L}|$, $W$ denotes the graph adjacency matrix $W_{uv}=w_{\alpha,\beta}(u,v)$, and $D$ is the diagonal matrix with entries $D_{ii}=\sum_{j}W_{ij}$. The final predictions are obtained by rounding $f_{u}$ for $u\in U$, i.e. predicting $\mathbb{I}\{f_{u}\geq\frac{1}{2}\}$, denoted by $G^{\sigma,\beta}(\mathcal{L},\mathcal{U},\bm{\delta})$. Let $v^{\sigma,\beta}:(\mathcal{L},\mathcal{U},\bm{\delta})\mapsto[0,1]$ denote the average 0-1 binary classification loss of the predictions of the above algorithm when the graph is built with parameters $\sigma,\beta$.
In the data-driven setting, we are given multiple problem instances $P_{1},\dots,P_{i}$, each $P_{i}$ is represented as a tuple $(\mathcal{L}_{i},\mathcal{U}_{i},\bm{\delta}_{i})$. Assuming that there is an underlying problem distribution, that represents a specific application domain, we want to know how many problems instance is sufficient to learn the best parameters $\alpha,\beta$ that is near-optimal for such problem distribution. To do that, we want to analyze the pseudo-dimension of the following function class:
$\mathcal{G}=\{v^{\sigma,\beta}:(\mathcal{L},\mathcal{U},\bm{\delta})\mapsto v(G^{\sigma,\beta}(\mathcal{L},\mathcal{U},\bm{\delta}))\mid\sigma\in\mathbb{R}\setminus\{0\},\beta\in\Delta([L])\}.$
6.2.2 Generalization guarantee for data-driven semi-supervised learning with Gaussian RBF kernel and multiple distance functions
We now instantiate the main result in this section, which establishes an upper bound for the pseudo-dimension of the function class $\mathcal{G}$.
Theorem 6.4.
Let $\mathcal{G}$ be a class of functions mapping semi-supervised learning instances $(\mathcal{L},\mathcal{U},\bm{\delta})$ to $[0,1]$. Then $\text{Pdim}(\mathcal{G})=\mathcal{O}(n^{4}L^{2})$, where $n=\left|\mathcal{L}\right|+\left|\mathcal{U}\right|$ is the total number of samples in each problem instance, and $L=\left|\bm{\delta}\right|$ is the number of distance functions.
Proof.
Fix a problem instance $(\mathcal{L},\mathcal{U},\bm{\delta})$, we will show that the dual loss function $v_{\mathcal{L},\mathcal{U},\bm{\delta}}(\sigma,\beta):=v^{\sigma,\beta}(\mathcal{L},\mathcal{U},\bm{\delta})$ is piece-wise constant and characterize the number and complexity of the boundary functions which we will show belong to a Pfaffian system. Our main result implies a bound on the pseudo-dimension of $\mathcal{G}$. The quadratic objective minimization has a closed-form solution [ZGL03], given by
$f_{\mathcal{U}}=(D_{\mathcal{U}\mathcal{U}}-W_{\mathcal{U}\mathcal{U}})^{-1}W_{\mathcal{U}\mathcal{L}}f_{\mathcal{L}},$
where subscripts $\mathcal{U},\mathcal{L}$ refer to the unlabeled and labeled points, respectively. Here $f_{\mathcal{L}}\in\{0,1\}^{|\mathcal{L}|}$ is the ground truth label of samples in the labeled set $\mathcal{L}$.
The key challenge now is to analyze the formula $(D_{\mathcal{U}\mathcal{U}}-W_{\mathcal{U}\mathcal{U}})^{-1}$ of which the element will be shown to be Pfaffian functions of $\sigma,\beta$. Recall that $w_{u,v}=\exp(-\delta_{\beta}(u,v)/\sigma^{2})$ for $u,v\in\mathcal{L}\cup\mathcal{U}$, and $D_{ii}=\sum_{j=1}^{n}w_{ij}$. First, we recall the identity $A^{-1}=\frac{\text{adj}(A)}{\det(A)}$, for any invertible matrix $A$, where $\text{adj}(A)$ and $\det{A}$ are the adjoint and determinant of matrix $A$, respectively. Therefore, we can see that any element of $(D_{\mathcal{U}\mathcal{U}}-W_{\mathcal{U}\mathcal{U}})^{-1}$ is a rational function of $w_{u,v}$ of degree at most $\left|\mathcal{U}\right|$.
Now, consider the Pfaffian chain $\mathcal{C}(\sigma,\beta,\frac{1}{\sigma},w_{11},\dots,w_{nn})$ with $L+1$ variables $\sigma,\beta$, and of length $q=n^{2}+1$. To see the Pfaffian degree of $\mathcal{C}$, note that for any pair of nodes $u,v\in\mathcal{U}\cup\mathcal{L}$, we have
$\frac{\partial w_{u,v}}{\partial\beta_{i}}=-\frac{\delta_{i}(u,v)}{\sigma^{2}}\exp\left(-\frac{\delta_{\beta}(u,v)}{\sigma^{2}}\right)=-\delta_{i}(u,v)g^{2}w_{u,v},$
$\frac{\partial w_{u,v}}{\partial\alpha}=\frac{2\delta_{\beta}(u,v)}{\sigma^{3}}\exp\left(-\frac{\delta_{\beta}(u,v)}{\sigma^{2}}\right)=2\delta_{\beta}(u,v)g^{3}w_{u,v}.$
Thus, the Pfaffian chain $\mathcal{C}$ has Pfaffian degree $M=5$.
Now, consider the dual loss function $v_{\mathcal{L},\mathcal{U},\bm{\delta}}(\sigma,\beta)$. Note that
$v_{\mathcal{L},\mathcal{U},\bm{\delta}}(\sigma,\beta)=\frac{1}{\left|\mathcal{U}\right|}\sum_{u\in\mathcal{U}}\mathbb{I}(\mathbb{I}(f_{u}\geq 1/2)=g_{u}),$
where $g_{u}$ is the ground-truth label of unlabeled node $u$, using for evaluation purpose. We can see that, for each $u\in\mathcal{U}$, $\mathbb{I}(f_{u}\geq 1/2)=\mathbb{I}\left(\frac{f^{(1)}_{u}}{f^{(2)}_{u}}\geq 1/2\right)=\mathbb{I}\left(f^{(1)}_{u}\geq 1/2f^{(2)}_{u}\right)$ is a boundary function. Both functions $f^{(1)}_{u},f^{(2)}_{u}$ are Pfaffian functions from chain $\mathcal{C}$ and of degree $|\mathcal{U}|$. In each region determined by the sign vector $b_{\mathcal{U}}$, where $b_{u}=\mathbb{I}(f_{u}\geq 1/2)$ for $u\in\mathcal{U}$, the dual loss function $v_{\mathcal{L},\mathcal{U},\bm{\delta}}(\sigma,\beta)$ is a constant functions. We can see that there are at most $\left|\mathcal{U}\right|+1$ such constant functions that $v_{\mathcal{L},\mathcal{U},\bm{\delta}}(\sigma,\beta)$ can take. Therefore, by Definition 8, the dual function class $\mathcal{G}^{*}$ is $(\left|\mathcal{U}\right|+1,\left|\mathcal{U}\right|+1,n^{2}+1,5,\left|\mathcal{U}\right|)$-Pfaffain piece-wise structured. We can apply Theorem 5.2 to get
$\text{Pdim}(\mathcal{G})=\mathcal{O}(n^{4}L^{2}+n^{2}L\log(|\mathcal{U}|)+n^{2}L\log L+L\log|\mathcal{U}|).$
Noting $|\mathcal{U}|\leq n$ and simplifying yields the claimed bound. ∎
The solution to the harmonic objective of Zhu et al. [ZGL03] has a closed-form expresion $f_{\mathcal{U}}=(D_{\mathcal{U}\mathcal{U}}-W_{\mathcal{U}\mathcal{U}})^{-1}W_{\mathcal{U}\mathcal{L}}f_{\mathcal{L}}$. Here each element of $(D_{\mathcal{U}\mathcal{U}}-W_{\mathcal{U}\mathcal{U}})^{-1}$ can be shown to be a Pfaffian function of the graph hyperparameter $\sigma$ over a Pfaffian chain of length $n^{2}+1$. We then apply Theorem 5.2 to obtain the stated pseudo-dimension bound.
7 Online Learning
In this section, we introduce new general techniques for establishing no-regret learning guarantee for data-driven algorithms in the online learning setting. Balcan et al. [BDP20] provided a tool for verifying dispersion when non-Lipschitzness (or discontinuity) occurs along roots of polynomials in one and two dimensions, and Balcan et al. [BS21] extended the result to algebraic varieties in higher dimensions. Here, we further generalize the results to cases where non-Lipschitzness occurs along Pfaffian hypersurfaces.
We first establish a constant upper bound on the VC-dimension of a class of Pfaffian functions with bounded chain length and Pfaffian degree when labeled by axis-aligned segments. This bound is instrumental in extending the framework for establishing dispersion developed in prior work [BDP20, BS21]. The necessary background from the theory of Pfaffian functions to comprehend these results is provided in Appendix E.
Theorem 7.1.
There is a constant $K_{M,q,d,p}$ depending only on $M$, $q$, $d$, and $p$ such that axis-aligned line segments in $\mathbb{R}^{p}$ cannot shatter any collection of $K_{M,q,d,p}$ Pfaffian hypersurfaces consisting of functions from a common chain of length $q$, degree at most $d$ and Pfaffian degree at most $M$.
Proof.
Let $\mathcal{C}(\bm{x},f_{1},\dots,f_{q})$, where $\bm{x}\in\mathbb{R}^{p}$, is a Pfaffian chain of length $q$ and of Pfaffian degree at most $M$. Let $\mathcal{P}=\{P_{1},\dots,P_{K}\}$ denote a collection of $K$ Pfaffian hypersurfaces $P_{i}(\bm{x},f_{1}(\bm{x}),\dots,f_{q}(\bm{x}))=0$ for $i=1,\dots,K$, where each $P_{i}$ is a Pfaffain function of degree at most $d$ from the Pfaffian chain $\mathcal{C}$. We say that a subset $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ is hit by a line segment $\upsilon$ if for any $P\in\mathcal{P}^{\prime}$, the hypersurface $P=0$ intersect with $\upsilon$. We say that $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ is hit by a line $\Upsilon$ if for any $P\in\mathcal{P}^{\prime}$, there exist a line segment $\upsilon\subset\Upsilon$ such that the hypersurface $P=0$ intersects with $\upsilon$. We seek to upper bound the number of subsets of $\mathcal{P}$ which may be hit by axis-aligned line segments. We will first consider shattering by line segments in a fixed axial direction $x=\bm{x}_{j}$ for $j\in[p]$.
Let $\Upsilon_{j}$ be a line along the axial direction $\bm{x}_{j}$. The subsets of $\mathcal{P}$ which may be hit by segments along $\Upsilon_{j}$ are determined by the pattern of intersections of $\Upsilon_{j}$ with the hypersurfaces in $\mathcal{P}$. We can now use Theorem E.1 to bound the number of intersections between $\Upsilon_{j}$ and any hypersurface $P_{i}=0$ for $P_{i}\in\mathcal{P}$ as
$R:=2^{q(q-1)/2}d(M\min\{q,p\}+d)^{q},$
using the fact that the straight line $\Upsilon_{j}$ is given by $p-1$ equations $\bm{x}_{k}=0$ for $k\neq j$ which form a Pfaffian system with $P_{i}$ of chain length $q$, Pfaffian degree $M$, and degrees $\deg(P_{i})=d$, $\deg(x_{k})=1$ for $k\neq j$. Therefore $\Upsilon_{j}$ intersects hypersurfaces in $\mathcal{P}$ at most $KR$ points, resulting in at most $KR+1$ segments of $\Upsilon_{j}$. Thus, $\Upsilon_{j}$ may hit at most (${KR+1\choose 2}=\mathcal{O}(K^{2}2^{q(q-1)}d^{2}{(M\min\{q,p\}+d)^{2q}})$ subsets of $\mathcal{P}$.
We will now bound the number of distinct subsets that may be hit as the axis-aligned $\Upsilon_{j}$ is varied (while it is still along the direction $\bm{x}_{j}$). Define the equivalence relation $L_{x_{1}}\sim L_{x_{2}}$ if the same sequence of hypersurfaces in $\mathcal{P}$ intersect $L_{x_{1}}$ and $L_{x_{2}}$ (in the same order, including with multiplicities). To obtain these equivalence classes, we will project the hypersurfaces in $\mathcal{P}$ onto a hyperplane orthogonal to the $\bm{x}_{j}$-direction. By the generalization of the Tarski-Seidenberg theorem to Pfaffian manifolds, we get a collection of semi-Pfaffian sets [vdD86], and these sets form a cell complex with at most $\mathcal{O}(K^{(p!)^{2}(2p(2p+q))^{p}}(M+d)^{q^{\mathcal{O}(p^{3})}})$ cells [GV01]. This is also a bound on the number of equivalence classes for the relation $\sim$.
Putting together, for each axis-parallel direction $\bm{x}_{j}$, the number of distinct subsets of $\mathcal{P}$ hit by any line segment along the direction $\bm{x}_{j}$ is at most $\mathcal{O}(K^{2}2^{q(q-1)}d^{2}{(M\min\{q,p\}+d)^{2q}}K^{(p!)^{2}(2p(2p+q))^{p}}(M+d)^{q^{\mathcal{O}(p^{3})}})$. Thus, for all axis-parallel directions, we get that the total number of subsets of $\mathcal{P}$ that may be hit is at most
$C_{K}=O\left(p2^{q(q-1)}d^{2}{(M\min\{q,p\}+d)^{2q}}K^{(p!)^{2}(2p(2p+q))^{p}+2}(M+d)^{q^{\mathcal{O}(p^{3})}}\right).$
For fixed $M,q,d,p$, this grows sub-exponentially in $K$, and therefore there is an absolute constant $K_{M,q,d,p}$ such that $C_{K}<2^{K}$ provided $K\geq K_{M,q,d,p}$. This implies that a collection of $K_{M,q,d,p}$ Pfaffian hypersurfaces cannot be shattered by axis-aligned line segments and establishes the desired claim. ∎
We can now use the above theorem to establish the following general tool for verifying dispersion of piece-wise-Lipschitz functions where the piece boundaries are given by Pfaffian hypersurfaces. A full proof of the following result is located in Appendix E.
Theorem 7.2.
Let $l_{1},\dots,l_{T}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ be independent piece-wise $L$-Lipschitz functions, each having discontinuities specified by a collection of at most $K$ Pfaffian hypersurfaces of bounded degree $d$, Pfaffian degree $M$ and length of common Pfaffian chain $q$. Let $L$ denote the set of axis-aligned paths between pairs of points in $\mathbb{R}^{p}$, and for each $s\in L$ define $D(T,s)=|\{1\leq t\leq T\mid l_{t}\text{ has a discontinuity along }s\}|$. Then we have that $\mathbb{E}[\sup_{s\in L}D(T,s)]\leq\sup_{s\in L}\mathbb{E}[D(T,s)]+\tilde{O}(\sqrt{T\log T})$, where the soft-O notation suppresses constants in $d,p,M,q$.
Proof Sketch.
We relate the number of ways line segments can label vectors of $K$ Pfaffian hypersurfaces of degree $d$, Pfaffian degree $M$ and common chain length $q$, to the VC-dimension of line segments (when labeling Pfaffian hypersurfaces), which from Theorem 7.1 is constant. To verify dispersion, we need a uniform-convergence bound on the number of Lipschitzness violations between the worst pair of points $\rho,\rho^{\prime}$ at distance $\leq\epsilon$, but the definition allows us to bound the worst rate of discontinuties along any path between $\rho,\rho^{\prime}$ of our choice. We can bound the VC-dimension of axis aligned segments against Pfaffian hypersurfaces of bounded complexity, which will allow us to establish dispersion by considering piece-wise axis-aligned paths between points $\rho$ and $\rho^{\prime}$. ∎
The above result reduces the problem of verifying dispersion to showing that the expected number of discontinuities $\mathbb{E}[D(T,s)]$ along any axis-aligned path $s$ is $\tilde{O}(\epsilon T)$, which together with Theorem 7.2 implies that the sequence of functions is $1/2$-dispersed.
7.1 Online learning for data-driven agglomerative hierarchical clustering
We illustrate how this can be used for learning the parameters online in agglomerative hierarchical clustering. We refer to Appendix E for proof details.
Theorem 7.3.
Consider an adversary choosing a sequence of clustering instances where the $t$-th instance has a symmetric distance matrix $D^{(t)}\in[0,R]^{n\times n}$ and for all $i\leq j,d^{(t)}_{ij}$ is $\kappa$-smooth. For the sub-family of $\mathcal{M}_{1}$ with $\alpha>\alpha_{\min}$ for $0<\alpha_{\min}<\infty$, we have that the corresponding sequence of utility functions $u_{1}^{(t)}$ are $1/2$-dispersed.
Proof Sketch. By using a generalization of the Descartes’ rule of signs, we first show that there is at most one positive real solution for the equation
$(d(a_{1},b_{1}))^{\alpha}+(d(a_{2},b_{2}))^{\alpha}=(d(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha}+(d(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha}.$
We then use $\kappa$ boundedness of the distances $d_{ij}$ to show that the probability of having a zero $\alpha^{*}\in I$ in any interval of width $I$ is at most $\tilde{O}(\epsilon)$. Using that we have at most $n^{8}$ such boundary functions, we can conclude that $\mathbb{E}[D(T,s)]=\tilde{O}(n^{8}\epsilon T)$ and use Theorem 7.2. $\square$
We define another family of linkage clustering algorithms that uses $L$ parameters $\alpha=(\alpha_{1},\dots,\alpha_{L})$ to interpolate several metrics $d_{1},d_{2},\dots,d_{L}$ as follows
$\mathcal{M}_{3}$ =$=\Bigg{\{}m^{3,\alpha}_{d}:(A,B)\mapsto\left(\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(d_{i}(a,b))^{\alpha_{i}}\right)^{1/\sum_{i}\alpha_{i}}\mid\alpha_{i}\in[\alpha_{\min},\alpha_{\max}]\Bigg{\}},$
where $0<\alpha_{\min}<\alpha_{\max}\leq 1$. For this family, we have the following result.
Theorem 7.4.
Consider an adversary choosing a sequence of clustering instances where the $t$-th instance has a symmetric distance matrix $D^{(t)}\in[0,R]^{n\times n}$ and for all $i\leq j,d^{(t)}_{ij}$ is $\kappa$-smooth. For the family of clustering algorithms $\mathcal{M}_{3}$, we have that the corresponding sequence of utility functions $u_{3}^{(t)}$ as a function of the parameter $\alpha=(\alpha_{1},\dots,\alpha_{L})$ are $1/2$-dispersed.
Proof Sketch. It is sufficient to show dispersion for each $\alpha_{i}$ keeping the remaining parameters fixed. This is because, the definition of dispersion allows us to consider intersections of discontinuities with axis-aligned paths. WLOG, assume $\alpha_{2},\dots,\alpha_{L}$ are fixed. The boundary functions are given by solutions of exponential equations in $\alpha_{1}$ of the form
$\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(d_{i}(a,b))^{\alpha_{i}}=\frac{1}{|A^{\prime}||B^{\prime}|}\sum_{a^{\prime}\in A^{\prime},b^{\prime}\in B^{\prime}}\Pi_{i\in[L]}(d_{i}(a^{\prime},b^{\prime}))^{\alpha_{i}},$
for $A,B,A^{\prime},B^{\prime}\subseteq S$. We can now use Theorem 7.2 together with Theorem 25 of [BS21] to conclude that the discontinuities of $u_{3}^{(t)}$ are $\frac{1}{2}$-dispersed. $\square$
7.2 Online learning for data-driven regularized logistic regression
We consider an online learning setting for regularized logistic regression from Section D.2. The problem instances $P_{t}=(X^{(t)},y^{(t)},X^{(t)}_{\text{val}},y^{(t)}_{\text{val}})$ are now presented online in a sequence of rounds $t=1,\dots,T$, and the online learner predicts the regularization coefficient $\lambda_{t}$ in each round. The regret of the learner is given by
$R_{T}=\mathbb{E}\sum_{t=1}^{T}h(\lambda_{t},P_{t})-h(\lambda^{*},P_{t}),$
where $\lambda^{*}=\operatorname*{arg\,min}_{\lambda\in[\lambda_{\min},\lambda_{\max}]}\mathbb{E}\sum_{t=1}^{T}h(\lambda,P_{t})$. Our main result is to show the existence of a no-regret learner in this setting.
Theorem 7.5.
Consider the online learning problem for tuning the logistic regression regularization coefficient $\lambda_{t}$ stated above. There exists an online learner with expected regret bounded by $\mathcal{O}(\sqrt{T\log[(\lambda_{\text{max}}-\lambda_{\text{min}})T]})$.
Proof Sketch. The key idea is to use an appropriate surrogate loss function which well approximates the true loss function, but is dispersed and therefore can be learned online. We then connect the regret of the learner with respect to the surrogate loss with its regret w.r.t. the true loss.
We consider an $\epsilon$-grid of $\lambda$ values. For each round $t$, to construct the surrogate loss function, we sample a uniformly random point from each interval $[\lambda_{\text{min}}+k\epsilon,\lambda_{\text{min}}+(k+1)\epsilon]$ and compute the surrogate loss $h^{(\epsilon)}_{\lambda}(P)$ at that point, and use a linear interpolation between successive points. By Theorem D.1, this has an error of at most $\mathcal{O}(\epsilon^{2})$, which implies the regret gap with the true loss is at most $\mathcal{O}(\epsilon^{2}T)$ when using the surrogate function.
We show that the surrogate function is $f$-point-dispersed (Definition 4, [BDP20]) with $f(T,r)=\frac{rT}{\epsilon}$. Using Theorem 5 of [BDP20], we get that there is an online learner with regret $\mathcal{O}(\sqrt{T\log((\lambda_{\text{max}}-\lambda_{\text{min}})/r)}+f(T,r)+Tr+\epsilon^{2}T)=\mathcal{O}(\sqrt{T\log((\lambda_{\text{max}}-\lambda_{\text{min}})/r)}+\frac{Tr}{\epsilon}+\epsilon^{2}T)$. Setting $\epsilon=T^{-1/4}$ and $r=T^{-3/4}$, we get the claimed regret upper bound. $\square$
8 Conclusion and future directions
In this work, we introduce the Pfaffian GJ framework for the data-driven distributional learning setting, providing learning guarantees for function classes whose computations involve the broad class of Pfaffian functions. Additionally, we propose a refined Pfaffian piece-wise structure for data-driven algorithm design problems and establish improved learning guarantees for problems admitting such a refined piece-wise structure. We apply our framework to a variety of previously studied data-driven algorithm design problems (including hierarchical clustering, graph-based semi-supervised learning and regularized logistic regression) under novel and more challenging settings, where known techniques do not yield concrete learning guarantees. By carefully analyzing the underlying Pfaffian structure for these problems, we leverage our proposed framework to establish novel statistical learning guarantees. We further study the data-driven online learning setting, where we introduce a new tool for verifying the dispersion property, applicable when the transition boundaries of the utility functions involve Pfaffian functions.
Appendix A Preliminaries
A.1 Classical generalization results
For completeness, along with the definition of pseudo-dimension (Definition 1), we now give a formal definition of the VC-dimension, which is a counterpart of pseudo-dimension when the function class is binary-valued.
Definition 9 (Shattering and VC-dimension).
Let $\mathcal{F}$ is a class of binary-valued functions that take input from $\mathcal{X}$. Let $S=\{x_{1},\dots,x_{N}\}$, we say that $S$ is shattered by $\mathcal{F}$ if
|$|\{(f(x_{1}),\dots,f(x_{N}))\mid f\in\mathcal{F}\}|=2^{N}.$
The VC-dimension of $\mathcal{F}$, denoted $\text{VCdim}(\mathcal{F})$, is the largest size of the set $S$ that can be shattered by $\mathcal{F}$.
Rademacher complexity [BM02] is another classical learning-theoretic complexity measure for obtaining data-dependent learning guarantees.
Definition 10 (Rademacher complexity, [BM02, Wai19]).
Let $\mathcal{F}$ be a real-valued function class which takes input from domain $\mathcal{X}$. Let $\mathcal{S}=\{x_{1},\dots,x_{n}\}\subset\mathcal{X}$ be a finite set of inputs. Then the empirical Rademacher complexity $\mathscr{R}_{\mathcal{S}}(\mathcal{F})$ of $\mathcal{F}$ with respect to $\mathcal{S}$ is defined as
^$\hat{\mathscr{R}}_{\mathcal{S}}(\mathcal{F})=\mathbb{E}_{\bm{\sigma}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\sigma_{i}f(x_{i})\right|\right].$
The Rademacher complexity of $\mathcal{F}$ for $n$ samples is then defined as
$\mathscr{R}_{n}(\mathcal{F})=\mathbb{E}_{\mathcal{S}}[\hat{\mathscr{R}}_{\mathcal{S}}(\mathcal{F})]=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{\bm{\sigma}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\sigma_{i}f(x_{i})\right|\right].$
The following classical lemma allows us to bound the Rademacher complexity of a real-valued function class via its pseudo-dimension.
Lemma A.1.
Let $\mathcal{F}$ be a real-valued function class and bounded by $H$. Then we have
$\mathscr{R}_{n}(\mathcal{F})\leq O\left(H\sqrt{\frac{\text{Pdim}(\mathcal{F})}{H}}\right).$
A.2 Uniform convergence and PAC-learnability
We now remind classical notions in learning theory, which formalizes the learnability of a function class.
Definition 11 (Uniform convergence).
Let $\mathcal{F}$ is a real-valued function class which takes value from $\mathcal{X}$, and $\mathcal{D}$ is a distribution over $\mathcal{X}$. If for any $\epsilon>0$, and any $\delta\in[0,1]$, there exists $N(\epsilon,\delta)$ s.t. w.p. at least $1-\delta$ over the draw of $n\geq N(\epsilon,\delta)$ samples $X_{1},\dots,X_{n}\sim\mathcal{D}$, we have
$\Delta_{n}=\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\mathbb{E}_{X\sim\mathcal{D}}[f(X)]\right|\leq\epsilon,$
then we say that (the empirical process of) $\mathcal{F}$ uniformly converges. The quantity $\Delta_{n}$ is called the empirical process of function class $\mathcal{F}$.
The following classical result demonstrates the connection between uniform convergence and learnability with an ERM learner.
Theorem A.2 ([MRT18]).
If $\mathcal{F}$ has a uniform convergence guarantee with $s(\epsilon,\delta)$ samples then it is PAC learnable with ERM and $s(\epsilon/2,\delta)$ samples.
Proof.
For $S=\{x_{1},\dots,x_{N}\}$, let $L_{S}(f)=\frac{1}{n}\sum_{i=1}^{n}f(x_{i})$, and $L_{\mathcal{D}}(f)=\mathbb{E}_{X\sim\mathcal{D}}[f(X)]$ for $f\in\mathcal{F}$. Since $\mathcal{F}$ is uniform convergence with $s(\epsilon,\delta)$ samples, w.p. at least $1-\delta$ for all $f\in\mathcal{F}$, we have $|L_{S}(f)-L_{\mathcal{D}}(f)|$ for any set $S$ with cardinal $m\geq s(\epsilon,\delta)$. Let $f_{ERM}\in\operatorname*{arg\,min}_{f\in\mathcal{F}}L_{S}(f)$ be the hypothesis outputted by the ERM learner, and $f^{*}\in\operatorname*{arg\,min}_{f\in\mathcal{F}}L_{\mathcal{D}}(f)$ be the best hypothesis. We have
$L_{\mathcal{D}}(f_{ERM})\leq L_{S}(f_{ERM})+\frac{\epsilon}{2}\leq L_{S}(f^{*})+\frac{\epsilon}{2}\leq L_{\mathcal{D}}(h^{*})+\epsilon,$
which concludes the proof. ∎
Appendix B Omitted proofs from Section 4
In this section, we will present basic properties of Pfaffian chain/functions and a proof for Theorem 4.2.
B.1 Some facts about Pfaffian functions with elementary operators
In this section, we formalize some basic properties of Pfaffian functions. Essentially, the following results say that if adding/subtracting/multiplying/dividing two Pfaffian functions from the same Pfaffian chain, we end up with a function that is also the Pfaffian function from that Pfaffian chain. Moreover, even if the two Pfaffian functions are not from the same Pfaffian chain, we can still construct a new Pfaffian chain that contains both the functions and the newly computed function as well (by simply combining the two chains).
Fact 1 (Addition/Subtraction).
Let $g$, $h$ be Pfaffian functions from the Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$ ($a\in\mathbb{R}^{d}$). Then we have $g(\bm{a})\pm h(\bm{a})$ are also Pfaffian functions from the chain $\mathcal{C}$.
Proof.
For any $a_{i}$, we have $\frac{\partial}{\partial a_{i}}(g\pm h)(\bm{a})=\frac{\partial}{\partial a_{i}}g(\bm{a})\pm\frac{\partial}{\partial a_{i}}h(\bm{a})$. Since $\frac{\partial g(\bm{a})}{\partial a_{i}}$ and $\frac{\partial h(\bm{a})}{\partial a_{i}}$ are polynomial of $\bm{a}$ and $\eta_{i}$, $\frac{\partial}{\partial a_{i}}(g\pm h)(\bm{a})$ are also polynomial of $\bm{a}$ and $\eta_{i}$, which concludes the proof. ∎
Fact 2 (Multiplication).
Let $g$, $h$ be Pfaffian functions from the Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$ ($\bm{a}\in\mathbb{R}^{d}$) of length $q$ and Pfaffain degree $M$. Then we have $g(\bm{a})h(\bm{a})$ is a Pfaffian function from the chain $\mathcal{C}^{\prime}(\bm{a},\eta_{1},\dots,\eta_{q},g,h).$
Proof.
For any $a_{i}$, we have $\frac{\partial}{\partial a_{i}}(g(\bm{a})h(\bm{a}))=g\frac{\partial h(\bm{a})}{\partial a_{i}}+h\frac{\partial g(\bm{a})}{\partial a_{i}}$, which is a polynomial of $\bm{a},\eta_{1},\dots,\eta_{q},g,h$. ∎
Fact 3 (Division).
Let $g,h$ be Pfaffain functions from the Pfaffian chain $\mathcal{C}(\bm{a},\eta_{q},\dots,\eta_{q})$ ($\bm{a}\in\mathbb{R}^{d}$) of length $q$ and Pfaffian degree $M$. Assume that $h(\bm{a})\neq 0$, then we have $\frac{g(\bm{a})}{h(\bm{a})}$ is a Pfaffian function from the Pfaffian chain $\mathcal{C}^{\prime}(\bm{a},\eta_{1},\dots,\eta_{q},,\frac{1}{h},\frac{g}{h})$.
Proof.
For any $a_{i}$, we have $\frac{\partial}{\partial a_{i}}\left(\frac{g(\bm{a})}{h(\bm{a})}\right)=\frac{\partial g(\bm{a})}{\partial a_{i}}\frac{1}{h(\bm{a})}-\frac{g(\bm{a})}{h(\bm{a})}\frac{\partial h(\bm{a})^{2}}{\partial a_{i}}$ is a polynomial of $\bm{a},\eta_{1},\dots,\eta_{q},\frac{1}{h},\frac{g}{h}$. ∎
Fact 4 (Composition).
Let $h$ be a Pfaffian function from the Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$ ($\bm{a}\in\mathbb{R}^{d}$), and $g$ be Pfaffian function from the Pfaffain chain $\mathcal{C}^{\prime}(b,\eta^{\prime}_{1},\dots,\eta^{\prime}_{q^{\prime}})$ ($b\in\mathbb{R}$). Then $g(h(\bm{a}))$ is a Pfaffian function from the Pfaffian chain $\mathcal{C}^{\prime\prime}(\bm{a},\eta_{1},\dots,\eta_{q},h,\eta^{\prime}_{1}(h),\dots,\eta^{\prime}_{q^{\prime}}(h)).$
Proof.
For any $a_{i}$, we have $\frac{\partial g(h(\bm{a}))}{\partial a_{i}}=\frac{\partial h(\bm{a})}{\partial a_{i}}\frac{\partial g(h(\bm{a}))}{\partial h(\bm{a})}$. Note that $\frac{\partial g(h(\bm{a}))}{\partial h(\bm{a})}=P(h(\bm{a}),\eta^{\prime}_{1}(h(a)),\dots,\eta^{\prime}_{q^{\prime}}(h(\bm{a}))$, where $P$ is some polynomial. Therefore, $g(h(\bm{a}))$ is a Pfaffian function from the Pfaffian chain
$\mathcal{C}^{\prime\prime}(\bm{a},\eta_{1},\dots,\eta_{q},h,\eta^{\prime}_{1}(h),\dots,\eta^{\prime}_{q^{\prime}}(h)).$
B.2 Proof for Theorem 4.2
In this section, we will present the proof of the Pfaffian GJ algorithm guarantee (Theorem 4.2). To begin with, we first recall some preliminary results about a standard technique for establishing pseudo-dimension upper-bound by analyzing the solution set connected components bound.
B.2.1 Preliminaries on the connection between pseudo-dimension and solution set connected components bound
We first introduce the notion of a regular value. Roughly speaking, given a smooth map $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, where $r\leq d$, a regular value $\epsilon\in\mathbb{R}^{r}$ is the point in the image space such that the solution of $F(a)=\epsilon$ is well-behaved.
Definition 12 (Regular value, [MW97]).
Consider $C^{\infty}$ functions $\Theta_{1},\dots,\Theta_{r}:\mathcal{A}\rightarrow\mathbb{R}$ where $r\leq d$, $\mathcal{A}\subseteq\mathbb{R}^{d}$, and the mapping $F:\mathcal{A}\rightarrow\mathbb{R}^{r}$ given by $F(\bm{a})=(\Theta_{1}(\bm{a}),\dots,\Theta_{r}(\bm{a}))$. Then $(\epsilon_{1},\dots,\epsilon_{r})\in\mathbb{R}^{r}$ is a regular value of F if either: (1) $F^{-1}((\epsilon_{1},\dots,\epsilon_{r}))=\emptyset$, or (2) $F^{-1}((\epsilon_{1},\dots,\epsilon_{r}))$ is a $(d-r)$-dimensional sub-manifold of $\mathbb{R}^{d}$.
It is widely known that by Sard’s Lemma [MW97], the set of non-regular values of the smooth map $F$ has Lebesgue measure 00. Based on this definition, we now present the definition of the solution set connected components bound.
Definition 13 (Solution set connected components bound, [KM97]).
Consider functions $\tau_{1},\dots,\tau_{K}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Given $\bm{x}_{1},\dots,\bm{x}_{N}$, assume that $\tau_{i}(\bm{x}_{j},\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a $C^{\infty}$ function for $i\in[K]$ and $j\in[N]$. For any $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, where $F(\bm{a})=(\Theta_{1}(\bm{a}),\dots,\Theta_{r}(\bm{a}))$ with $\Theta_{i}$ chosen from the $NK$ functions $\tau_{i}(\bm{x}_{j})$, if the number of connected components of $F^{-1}(\epsilon)$, where $\epsilon$ is a regular value, is upper bounded by $B$ independently on $x_{i}$, $r$, and $\epsilon$, then we say that $B$ is the solution set connected components bound.
The solution set connected components bound offers an alternative way of analyzing VC-dimension (or pseudo-dimension) of related parameterized function classes, which is formalized in the following lemma [KM97]. We include a high-level proof sketch for comprehensiveness.
Lemma B.1 ([KM97]).
Consider the binary-valued function $\Phi(\bm{x},\bm{a})$, for $\bm{x}\in\mathcal{X}$ and $\bm{a}\in\mathbb{R}^{d}$ constructed using the boolean operators AND and OR, and boolean predicates in one of the two forms ``$\tau(\bm{x},\bm{a})>0"$ or ``$\tau(\bm{x},\bm{a})=0"$. Assume that the function $\tau(\bm{x},\bm{a})$ can be one of at most $K$ forms ($\tau_{1},\dots,\tau_{K}$), where $\tau_{i}(\bm{x},\cdot)$ ($i\in[K])$ is a $C^{\infty}$ function of $\bm{a}$ for any fixed $x$. Let $\mathscr{C}_{\Phi}=\{\Phi_{\bm{a}}:\mathcal{X}\rightarrow\{0,1\}\mid\bm{x}\in\mathbb{R}^{d}\}$ where $\Phi_{\bm{a}}=\Phi(\cdot,\bm{a})$. Then
$\text{VCdim}(\mathscr{C}_{\Phi})\leq 2\log_{2}B+(16+2\log K)d,$
where $B$ is the solution set connected components bound.
Proof Sketch. The key idea involved in proving the following lemma is a combinatorial argument due to Warren (Theorem 1, [War68]) which bounds the number of connected components induced by a collection of boundary functions by $\sum_{j=0}^{d}b_{j}$, where $b_{j}$ is the number of connected components induced by intersections of any $j$ functions. This can be combined with the solution set connected components bound $B$, to get a bound $\sum_{j=0}^{d}2^{j}{NK\choose j}B\leq B\left(\frac{2NKe}{d}\right)^{d}$ on the total number of connected components. The result follows from noting $2^{N}\leq B\left(\frac{2NKe}{d}\right)^{d}$ if the $N$ instances $\bm{x}_{1},\dots,\bm{x}_{N}$ are to be shattered. $\square$
We now recall a result which is especially useful for bounding the solution set connected components bound $B$ for equations related to Pfaffian functions.
Lemma B.2 ([Kho91]).
Let $\mathcal{C}$ be a Pfaffian chain of length $q$ and Pfaffian degree $M$, consists of functions $f_{1},\dots,f_{q}$ in $a\in\mathbb{R}^{d}$. Consider a non-singular system of equations $\Theta_{1}(\bm{a})=\dots=\Theta_{r}(\bm{a})=0$ where $r\leq d$, in which $\Theta_{i}(\bm{a})$ ($i\in[r]$) is a polynomial of degree at most $\Delta$ in the variable $\bm{a}$ and in the Pfaffian functions $f_{1},\dots,f_{q}$. Then the manifold of dimension $k=d-r$ determined by this system has at most $2^{q(q-1)}\Delta^{m}S^{d-r}[(r-d+1)S-(r-d)]^{q}$ connected components, where $S=r(\Delta-1)+dM+1$.
The following corollary is the direct consequence of Lemma B.2.
Corollary B.3.
Consider the setting as in Lemma B.1. Assume that for any fixed $\bm{x}$, $\tau_{i}(\bm{x},\cdot)$ ($i\in[K]$) is a Pfaffian function from a Pfaffian chain $\mathcal{C}$ with length $q$ and Pfaffain degree $M$. Then
$B\leq 2^{dq(dq-1)/2}\Delta^{d}[(d^{2}(\Delta+M)]^{dq}.$
The following lemma gives a connection between the number of sign patterns and number of connected components.
Lemma B.4 (Section 1.8, [War68]).
Given $N$ real-valued functions $h_{1},\dots,h_{N}$, the number of distinct sign patterns |$\left|\{({\text{sign}}(h_{1}(\bm{a})),\dots,{\text{sign}}(h_{N}(\bm{a})))\mid\bm{a}\in\mathbb{R}^{d}\}\right|$ is upper-bounded by the number of connected components $\mathbb{R}^{d}-\cup_{i\in[N]}\{\bm{a}\in\mathbb{R}^{d}\mid g_{i}(\bm{a})=0\}$.
The following result is about the relation between the connected components and the solution set connected components.
Lemma B.5 (Lemma 7.9, [AB09]).
Let $\{f_{1},\dots,f_{N}\}$ be a set of differentiable functions that map from $\mathbb{R}^{d}\rightarrow\mathbb{R}$, with regular zero-set intersections. For each $i$, define $Z_{i}$ the zero-set of $f_{i}$: $Z_{i}=\{\bm{a}\in\mathbb{R}^{d}:f_{i}(\bm{a})=0\}$. Then
$CC\left(\mathbb{R}^{d}-\bigcup_{i\in[N]}Z_{i}\right)\leq\sum_{S\subseteq[N]}CC\left(\bigcap_{i\in S}Z_{i}\right).$
Combining with Definition 13, the RHS becomes $B\sum_{i=0}^{d}{N\choose i}\leq B\left(\frac{eN}{d}\right)^{d}$ for $N\geq d$.
B.2.2 Pfaffian formulae
We now introduce the building block of the Pfaffian GJ framework, named Pfaffian formula. Roughly speaking, a Pfaffian formula is a boolean formula that incorporates Pfaffian functions.
Definition 14 (Pfaffian formulae).
A Pfaffian formulae $f:\mathbb{R}^{d}\rightarrow\{True,False\}$ is a disjunctive normal form (DNF) formula over boolean predicates of the form $g(\bm{a},\eta_{1},\dots,\eta_{q})\geq 0$, where $g$ is a Pfaffian function of a Pfaffian chain $\mathcal{C}(\bm{a},\eta_{1},\dots,\eta_{q})$.
The following lemma essentially claims that for any function class, if its computation can be described by a Pfaffian formula and the corresponding Pfaffian chain exhibits bounded complexities, then the function class also possesses bounded pseudo-dimension.
Lemma B.6.
Suppose that each algorithm $L\in\mathcal{L}$ is parameterized by $\bm{a}\in\mathbb{R}^{d}$. Suppose that for every $\bm{x}\in\mathcal{X}$ and $r\in\mathbb{R}$, there is a Pfaffian formula $f_{\bm{x},r}$ of a Pfaffian chain $\mathcal{C}$ with length $q$ and Pfaffian degree $M$, that given $L\in\mathcal{L}$, check whether $L(\bm{x})>r$. Suppose that $f_{\bm{x},r}$ has at most $K$ distinct Pfaffian functions in its predicates, each of degree at most $\Delta$. Then,
$\text{Pdim}(\mathcal{L})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta K+16d.$
Proof.
This lemma is a direct consequence of Lemma B.1, and Corollary B.3. ∎
B.2.3 Proof for Theorem 4.2
We are now ready to present the proof of Theorem 4.2.
Theorem 4.2 (restated).
Consider a real-valued function class $\mathcal{L}$ with domain $\mathcal{X}$, of which each algorithm $L_{\bm{a}}\in\mathcal{L}$ is parameterized by ${\bm{a}}\in\mathcal{A}\subseteq\mathbb{R}^{d}$. Suppose that for every $\bm{x}\in\mathcal{X}$ and $r\in\mathbb{R}$, there is a Pfaffian GJ algorithm $\Gamma_{\bm{x},r}$, with associated Pfaffian chain $\mathcal{C}_{\bm{x},r}$ of length at most $q$ and Pfaffian degree at most $M$, that given $L_{\bm{a}}\in\mathcal{L}$, check whether $L_{\bm{a}}(\bm{x})>r$. Moreover, assume that values computed at intermediate steps of $\Gamma_{\bm{x},r}$ are from the Pfaffian chain $\mathcal{C}_{\bm{x},r}$, each of degree at most $\Delta$; and the functions computed in the conditional statements are of at most $K$ Pfaffian functions. Then $\text{Pdim}(\mathcal{L})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta K+16d.$
Proof.
We can transforms the computation of $T_{x,r}$ to a Pfaffian formula by converting the conditional statements into boolean combinations of Pfaffian formulae
$A\leftarrow\text{if $v\geq 0:X\text{ else }Y$ }\Rightarrow A\leftarrow((v\geq 0)\wedge X)\vee(\neg(v\geq 0)\wedge Y)$
where $v$, $X$, and $Y$ are values computed by Pfaffian formulae. Thus, we can apply Lemma B.6 to obtain the desired result $\text{Pdim}(\mathcal{L})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta K+16d.$
Appendix C Additional details and omitted proof for Section 5
C.1 Omitted proof for Section 5.2
Lemma 5.2 (restated).
Consider the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow[0,H]\mid{\bm{a}}\in\mathcal{A}\}$ where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Suppose that the dual function class $\mathcal{U}^{*}=\{u^{*}_{\bm{x}}:\mathcal{A}\rightarrow\mathbb{R}\mid\bm{x}\in\mathcal{X}\}$ is $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$-Pfaffian piece-wise structured, then the pseudo-dimension of $\mathcal{U}$ is bounded as follows
$\text{Pdim}(\mathcal{U})\leq d^{2}q^{2}+2dq\log(\Delta+M)+4dq\log d+2d\log\Delta(k_{\mathcal{F}}+k_{\mathcal{G}})+16d.$
Proof.
An intuitive explanation of this theorem can be found at Figure 2. Since $\mathcal{U}^{*}$ admits $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$-Pfaffian piece-wise structure, then for any problem instance $\bm{x}\in\mathcal{X}$ corresponding to the dual utility function $u^{*}_{\bm{x}}$, there are at most $\mathbb{I}(g_{\bm{x}}^{(1)}\geq 0),\dots,\mathbb{I}(g_{\bm{x}}^{(k)}\geq 0)$ where $k\leq k_{\mathcal{G}}$, and a piece function $f_{\bm{x},\bm{b}}$ for each binary vector $\bm{b}\in\{0,1\}^{k}$ such that for all $\bm{a}\in\mathcal{A}$, $u^{*}_{\bm{x}}(\bm{a})=f_{\bm{x},\bm{b}_{\bm{a}}}(\bm{a})$, where $\bm{b}_{\bm{a}}=(\mathbb{I}(g_{\bm{x}}^{(1)}(\bm{a})\geq 0),\dots,\mathbb{I}(g_{\bm{x}}^{(k)}(\bm{a})\geq 0))$. Therefore, for any problem instance $\bm{x}$ and real threshold $r\in\mathbb{R}$, the boolean value of $u^{*}_{\bm{x}}(\bm{a})-r\geq 0$ for any $\bm{a}\in\mathcal{A}$ can be calculated an algorithm $\Gamma_{\bm{x},r}$ described as follow: first calculating the boolean vector $\bm{b}_{\bm{a}}$, and then calculate the boolean value $f_{\bm{x},\bm{b}_{\bm{a}}}(\bm{a})-r\geq 0$.
From assumption, note that $g_{\bm{x}}^{(i)}$ and $f_{\bm{x},\bm{b}}$ are a Pfaffian functions of degree at most $\Delta$ from Pfaffian chain $\mathcal{C}_{\bm{x}}$ of length at most $q$ and Pfaffian degree at most $M$. This means that $\Gamma_{\bm{x},r}$ is a Pfaffian GJ algorithm. Therefore, combining with Theorem 4.2, we have the above claim. ∎
C.2 Omitted proofs for Section 5.3
Lemma 5.3 (restated).
Consider a function class $\mathcal{V}=\{v_{\bm{a}}:\mathcal{X}\rightarrow\mathbb{R}\mid{\bm{a}}\in\mathcal{A}\}$ where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Assume there is a partition $\mathcal{P}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{n}\}$ of the parameter space $\mathcal{A}$ such that for any problem instance $\bm{x}\in\mathcal{X}$, the dual utility function $u^{*}_{x}$ is a Pfaffian function of degree at most $\Delta$ in region $\mathcal{A}_{i}$ from a Pfaffian chain $\mathcal{C}_{\mathcal{A}_{i}}$ of length at most $q$ and Pfaffian degree $M$. Then the pseudo-dimension of $\mathcal{V}$ is upper-bounded as follows
$\text{Pdim}(\mathcal{V})=O(q^{2}d^{2}+qd\log(\Delta+M)+qd\log d+\log n).$
Proof.
Consider $N$ problem instances $\bm{x}_{1},\dots,\bm{x}_{N}$ with $N$ corresponding thresholds $\tau_{1},\dots,\tau_{N}$, we first want to bound the number of distinct sign patterns $\Pi_{\mathcal{V}}(N)=|\{({\text{sign}}(v_{\bm{a}}(\bm{x}_{1})-\tau_{1}),\dots,{\text{sign}}(v_{\bm{a}}(\bm{x}_{N})-\tau_{N}))\mid{\bm{a}}\in\mathcal{A}\}|$.
Denote $\Pi^{\mathcal{A}_{i}}_{\mathcal{V}}(N)=|\{({\text{sign}}(v_{\bm{a}}(\bm{x}_{1})-\tau_{1}),\dots,{\text{sign}}(v_{\bm{a}}(\bm{x}_{N})-\tau_{N}))\mid{\bm{a}}\in\mathcal{A}_{i}\}|$, we have $\Pi_{\mathcal{V}}(N)\leq\sum_{i=1}^{n}\Pi^{\mathcal{A}_{i}}_{\mathcal{V}}(N)$. For any $i\in[n]$, from the assumptions, Lemma 5.2 and using Sauer’s Lemma, we have
$\Pi_{\mathcal{V}}(N)\leq\left(\frac{eN}{S}\right)^{S},$
where $S=C(q^{2}d^{2}+qd\log(\Delta+M)+qd\log d)$ for some constant $C$. Therefore $\Pi_{\mathcal{V}}(N)\leq n\left(\frac{eN}{S}\right)^{S}$. Solving the inequality $2^{N}\leq\Pi_{\mathcal{V}}(N)$, we conclude that $\text{Pdim}(\mathcal{V})=O(q^{2}d^{2}+qd\log(\Delta+M)+qd\log d+\log n)$ as expected. ∎
C.3 A detailed comparison between the piece-wise structure by Balcan et al. [BDD+21] and our refined Pfaffian piece-wise structure
In this section, we will do a detailed comparison between our refined Pfaffian piece-wise structure, and the piece-wise structure proposed by Balcan et al. [BDD+21]. First, we derive concrete pseudo-dimension bounds implied by [BDD+21] when the refined Pfaffian piece-wise structure is present. We note that this implication is not immediate and needs a careful argument to bound the learning-theoretic complexity of Pfaffian piece and boundary functions which appear in their bounds.
Theorem C.1.
Consider the utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow\mathbb{R}\mid\bm{a}\in\mathcal{A}\}$, where $\mathcal{A}\subseteq\mathbb{R}^{d}$. Suppose that $\mathcal{U}^{*}$ admits $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$-Pfaffian piece-wise structure. Then using Theorem 5.1 by Balcan et al. [BDD+21], we have
$\text{Pdim}(\mathcal{U})=\mathcal{O}((d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d)\cdot\log[(d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d)k_{\mathcal{G}}]).$
Proof.
Consider an utility function class $\mathcal{U}=\{u_{\bm{a}}:\mathcal{X}\rightarrow\mathbb{R}\mid\bm{a}\in\mathcal{A}\}$, where $\mathcal{A}\subseteq\mathbb{R}^{d}$, we assume that $\mathcal{U}^{*}$ admits $(k_{\mathcal{F}},k_{\mathcal{G}},q,M,\Delta,d)$-Pfaffian piece-wise structure. Then, $\mathcal{U}^{*}$ also admits $(\mathcal{F},\mathcal{G},k_{\mathcal{G}})$ piece-wise structure following Definition 7. Here, $\mathcal{H}$ is the set of Pfaffian functions of degree $\Delta$ from a Pfaffian chain of length $q$ and Pfaffian degree $M$, and $\mathcal{G}$ is the set of threshold functions which are also Pfaffian functions of degree $\Delta$ from the same Pfaffian chain of length $q$ and Pfaffian degree $M$.
The first challenge in using the framework of Balcan et al. is that it only reduces the problem of bounding the learning-theoretic complexity of the piecewise-structured utility function to that of bounding the complexity of the piece and boundary functions involved, which is non-trivial in the case of Pfaffian functions. That is, we still need to bound $\text{Pdim}(\mathcal{F}^{*})$ and $\text{VCdim}(\mathcal{G}^{*})$, where $\mathcal{F}^{*}$ and $\mathcal{G}^{*}$ are dual function classes of $\mathcal{F}$ and $\mathcal{G}$, respectively. To bound $\text{VCdim}(\mathcal{G}^{*})$, we first consider the set of $N$ input instances $S=\{g_{1},\dots,g_{N}\}\subset\mathcal{G}$, and bound the number of distinct sign patterns,
$\Gamma_{S}(N)=|\{(g^{*}_{\bm{a}}(g_{1}),\dots,g^{*}_{\bm{a}}(g_{N}))\mid\bm{a}\in\mathcal{A}\}|=|\{(g_{1}(\bm{a}),\dots,g_{N}(\bm{a}))\mid\bm{a}\in\mathcal{A}\}|.$
From Lemma B.5 and B.4, and B.2, we can bound $\Gamma_{S}(N)$ as
$\Gamma_{S}(N)=\mathcal{O}(2^{dq(dq-1)/2}\Delta^{d}[(d^{2}(\Delta+M)])^{dq}\left(\frac{eN}{d}\right)^{d}.$
Solving the inequality $2^{N}=\mathcal{O}(2^{dq(dq-1)/2}\Delta^{d}[(d^{2}(\Delta+M)]^{dq}\left(\frac{ek}{d}\right)^{d})$, we have $\text{VCdim}(\mathcal{G}^{*})\leq\frac{dq(dq-1)}{2}+d\log\Delta+dq\log(d^{2}(\Delta+M))$.
We now bound $\text{Pdim}(\mathcal{F}^{*})$. By definition, we have =$\mathcal{F}^{*}=\{f^{*}_{\bm{a}}:\mathcal{F}\rightarrow\mathbb{R}\mid\bm{a}\in\mathcal{A}\}$. Again, to bound $\text{Pdim}(\mathcal{F}^{*})$, we first consider the set $S=\{f_{1},\dots,f_{N}\}\subset\mathcal{F}$ and a set of thresholds $T=\{r_{1},\dots,r_{N}\}$, and we want to bound the number of distinct sign patterns
$\Gamma_{S,T}(N)$ =|$=|\{({\text{sign}}(f^{*}_{\bm{a}}(f_{1})-r_{1}),\dots,{\text{sign}}(f^{*}_{\bm{a}}(f_{N})-r_{N}))\mid\bm{a}\in\mathcal{A}\}|$
=|$=|\{({\text{sign}}(f_{1}(\bm{a})-r_{1}),\dots,{\text{sign}}(f_{N}(\bm{a})-r_{N}))\mid\bm{a}\in\mathcal{A}\}|.$
Using similar argument as for $\mathcal{G}^{*}$, we have $\text{Pdim}(\mathcal{F}^{*})=\mathcal{O}(d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d)$. Combining with Theorem 3.3 by Balcan et al. [BDD+21], we conclude that
$\text{Pdim}(\mathcal{U})=\mathcal{O}((d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d)\log[(d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d)k_{\mathcal{G}}]).$
Remark 5.
Compared to our bounds in Theorem 5.2, Theorem C.1 (implied by [BDD+21]) has two notable differences. First, our bound is sharper by a logarithmic factor $\mathcal{O}(\log(d^{2}q^{2}+dq\log(\Delta+M)+dq\log d+d))$, which is a consequence of using a sharper form of the Sauer-Shelah lemma. Second, we have a dependence on a logarithmic term $k_{\mathcal{F}}$ which is asymptotically dominated by the other terms, and corresponds to better multiplicative constants than Theorem C.1.
Appendix D Additional results and omitted proofs for Section 6
We include complete proofs for upper bounds on the learning-theoretic complexity of linkage clustering algorithm families from Section 6.1. In Section D.2, we will leverage our proposed framework 5.3 to recover a recent result by Balcan et al. [BNS24].
D.1 Omitted proofs for Section 6.1
Theorem 6.2.
Let $\mathcal{H}_{2}$ be a class of functions
$\mathcal{H}_{2}=\{u^{\alpha,\beta}_{2}\colon(S,\bm{\delta})\mapsto u(A^{\alpha,\beta}_{2}(S,\bm{\delta}))\mid\alpha\in\mathbb{R}\cup\{-\infty,+\infty\},\beta\in\Delta([L])\}$
mapping clustering instances $(S,\bm{\delta})$ to $[0,1]$ by using merge functions from class $\mathcal{M}_{2}$ and an arbitrary merge function. Then $\text{Pdim}(\mathcal{H}_{2})=\mathcal{O}(n^{4}L^{2})$.
Proof.
Fix the clustering instance $(S,\bm{\delta})$. Suppose $A,B\subseteq S$ and $A^{\prime},B^{\prime}\subseteq S$ are two candidate clusters at some merge step of the algorithm. Then $A,B$ is preferred for merging over $A^{\prime},B^{\prime}$ iff
$\left(\frac{1}{|A||B|}\sum_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}\right)^{1/\alpha}\leq\left(\frac{1}{|A^{\prime}||B^{\prime}|}\sum_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}\right)^{1/\alpha},$
or equivalently,
$\frac{1}{|A||B|}\sum_{a\in A,b\in B}(\delta_{\beta}(a,b))^{\alpha}\leq\frac{1}{|A^{\prime}||B^{\prime}|}\sum_{a\in A^{\prime},b\in B^{\prime}}(\delta_{\beta}(a,b))^{\alpha}.$
For distinct choices of the point sets $A,B,A^{\prime},B^{\prime}$, we get at most (${2^{n}\choose 2}^{2}\leq 2^{4n}$ distinct boundary conditions across which the merge decision at any step of the algorithm may change.
We next show that the boundary functions constitute a Pfaffian system in $\alpha,\beta_{1},\dots,\beta_{L}$ and bound its complexity. For each pair of points $a,b\in S$, define $f_{a,b}(\alpha,\beta):=\frac{1}{\delta_{\beta}(a,b)}$, $g_{a,b}(\alpha,\beta):=\ln{\delta_{\beta}(a,b)}$ and $h_{a,b}(\alpha,\beta):={\delta_{\beta}(a,b)}^{\alpha}$. Similar to the proof of Theorem 6.1, these functions form a Pfaffian chain $\mathcal{C}(\alpha,\beta,f_{a,b},g_{a,b},h_{a,b})$ for $a,b\in S$. We can see that $\mathcal{C}$ is of length $q=3n^{2}$ and Pfaffian degree $M=2$. The boundary conditions can all be written in terms of the functions $\{h_{a,b}\}_{a,b\in S}$, meaning $k_{\mathcal{G}}=n^{8}$ and degree $\Delta=1$. Again, note that $k_{\mathcal{F}}=n+1$. Therefore, $\mathcal{H}_{2}^{*}$ admits $(n+1,n^{8},3n^{2},2,1,L+1)$-Pfaffian piece-wise structure. Applying Theorem 5.2 now gives that $\text{Pdim}(\mathcal{H}_{2})=\mathcal{O}(n^{4}L^{2}+n^{2}L\log L+L\log 2^{4n})=\mathcal{O}(n^{4}L^{2})$. ∎
Theorem 6.3.
Let $\mathcal{H}_{3}$ be a class of functions
$\mathcal{H}_{3}=\{u^{\alpha}_{3}:(S,\bm{\delta})\mapsto u(A^{\alpha}_{3}(S,\bm{\delta}))\mid\alpha_{i}\in\mathbb{R}\cup\{-\infty,\infty\}\setminus\{0\}\}$
mapping clustering instances $(S,\bm{\delta})$ to $[0,1]$ by using merge functions from class $\mathcal{M}_{3}$. Then $\text{Pdim}(\mathcal{H}_{3})=\mathcal{O}(n^{4}L^{2})$.
Proof.
Fix the clustering instance $(S,\bm{\delta})$. Suppose $A,B\subseteq S$ and $A^{\prime},B^{\prime}\subseteq S$ are two candidate clusters at some merge step of the algorithm. Then $A,B$ is preferred for merging over $A^{\prime},B^{\prime}$ iff
$\left(\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}\right)^{1/\sum_{i}\alpha_{i}}\leq\left(\frac{1}{|A^{\prime}||B^{\prime}|}\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}\right)^{1/\sum_{i}\alpha_{i}},$
or equivalently,
$\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}\leq\frac{1}{|A^{\prime}||B^{\prime}|}\sum_{a\in A^{\prime},b\in B^{\prime}}\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}.$
For distinct choices of the point sets $A,B,A^{\prime},B^{\prime}$, we get at most (${2^{n}\choose 2}^{2}\leq 2^{4n}$ distinct boundary conditions across which the merge decision at any step of the algorithm may change.
We next show that the boundary functions constitute a Pfaffian system in $\alpha_{1},\dots,\alpha_{L}$ and bound its complexity. For each pair of points $a,b\in S$, define $h_{a,b}(\alpha_{1},\dots,\alpha_{L}):=\Pi_{i\in[L]}(\delta_{i}(a,b))^{\alpha_{i}}$. Note that $\frac{\partial h}{\partial\alpha_{i}}:=\ln\delta_{i}(a,b)h_{a,b}(\alpha_{1},\dots,\alpha_{L})$, and thus these functions form a Pfaffian chain of chain length $n^{2}$ and Pfaffian degree 1. The boundary conditions can be all written in terms of the functions $\{h_{a,b}\}_{a,b\in S}$, meaning that $k_{\mathcal{G}}=2^{4n}$, and $\Delta=1$. Again, note that $k_{\mathcal{F}}=n+1$. Therefore $\mathcal{H}^{*}_{3}$ admits $(n+1,2^{4n},n^{2},1,1,L)$-Pfaffian piece-wise structure. Applying Theorem B.6 now gives that $\text{Pdim}(\mathcal{H}_{2})=\mathcal{O}(n^{4}L^{2})$. ∎
D.2 Additional result: Data-driven regularized logistic regression
Logistic regression [JWH+13] is a fundamental statistical technique and widely utilized classification model with numerous applications across diverse domains, including healthcare screening [ABM08], market forecasting [LB11], and engineering safety control [PD09], among others. To mitigate overfitting and enhance robustness, regularizations, such as sparsity ($\ell_{1}$) and shrinkage ($\ell_{2}$) constraints, are commonly employed in logistic regression [MRT18, Mur12]. In regularized logistic regression and regularized linear models in general, the regularization parameters, which control the magnitude of regularization, play a crucial role [MRT18, Mur12, JWH+13] and must be carefully chosen. Excessively high regularization parameters may lead to underfitting, while setting them too low may nullify the effectiveness of regularization. In practice, a common strategy for selecting regularization parameters is cross-validation, which is known to lack theoretical guarantees, in general, [Zha09, CLW16].
Recently, Balcan et al. [BNS24] proposed a data-driven approach for selecting regularization parameters in regularized logistic regression. Their methodology considers each regression problem, comprising training and validation sets, as a problem instance drawn from an underlying problem distribution. The objective is to leverage the available problem instances to determine regularization parameters that minimize the validation loss for any future problem instance sampled from the same distribution. In this section, we will leverage our proposed refined piece-wise structure to recover learning guarantees for the problem of data-driven logistic regression considered by [BNS24].
D.2.1 Problem setting
We recall the data-driven regularized logistic regression setting by Balcan et al. [BNS24]. Given a problem instance $P=(X,y,X_{\text{val}},y_{\text{val}})\in\mathcal{R}_{m,p,m^{\prime}}=\mathbb{R}^{m\times p}\times\{\pm 1\}^{m}\times\mathbb{R}^{m^{\prime}\times p}\times\{\pm 1\}^{m^{\prime}}$ where $(X,y)$ and $(X_{\text{val}},y_{\text{val}})$ are the training and validation set, respectively, and a regularization parameter $\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]$, the regularized logistic regression solves for the coefficients $\hat{\beta}_{(X,y)}$ that is the best fit for the training set $(X,y)$
$\hat{\beta}_{(X,y)}(\lambda)=\arg\min_{\beta\in\mathbb{R}^{p}}l(\beta,(X,y))+\lambda R(\beta),$
where $l(\beta,(X,y))=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(-y_{i}x_{i}^{\top}\beta))$ is the logistic loss, and $R(\beta)$ is either $\ell_{1}$ ($\|\beta\_1|$) or $\ell_{2}$($\|\beta\|^{2}_{2}$). After that, $\hat{\beta}_{(X,y)}(\lambda)$ is evaluated by $h(\lambda,P)$ which is the validation loss
$h(\lambda,P)=l(\hat{\beta}_{(X,y)}(\lambda),(X_{\text{val}},y_{\text{val}})).$
In the data-driven setting, we assume that there is an unknown problem distribution $\mathcal{D}$ over $\mathcal{R}_{m,p,m^{\prime}}$ that the problem instances are drawn from. Intuitively, the problem distribution $\mathcal{D}$ represents the application domain that the regularized logistic regression operates on. Our goal is to learn the best $\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]$ s.t. it minimizes the expected validation loss $\mathbb{E}_{P\sim\mathcal{D}}[h(\lambda,P)]$.
However, the problem distribution $\mathcal{D}$ is unknown. So instead we want to minimize $\mathbb{E}_{P\sim\mathcal{D}}[h(\lambda,P)]$ using ERM: assume that we are given $N$ problem instances $P^{(1)},\dots,P^{(N)}$ and we minimize the empirical validation loss
$\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]}\sum_{i=1}^{m}h(\lambda,P^{(i)}).$
The question now is: how many problem instances do we need to learn a good parameter $\hat{\lambda}$ so that the performance of $\mathbb{E}_{P\sim\mathcal{D}}[h(\hat{\lambda},P)$ is comparable to the performance of best parameter $\mathbb{E}_{P\sim\mathcal{D}}[h(\lambda^{*},P)]$, where $\lambda^{*}=\operatorname*{arg\,min}_{\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}]}}\mathbb{E}_{P\sim\mathcal{D}}[h(\lambda,P)]$. Define =$\mathcal{H}=\{h_{\lambda}:\mathcal{R}_{m,p,m^{\prime}}\rightarrow\mathbb{R}_{+}\mid\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]\}$. The goal above is equivalent to giving a learning guarantee for the function class $\mathcal{H}$.
D.2.2 Generalization guarantee for data-driven regularized logistic regression
Directly analyzing the learnability of $\mathcal{H}$ is challenging due to the unknown explicit form of $h(\lambda,P)$ as a function of $\lambda$. Instead, Balcan et al. [BNS24] propose analyzing the function class $\mathcal{H}$ via a surrogate validation loss function class $\mathcal{H}^{(\epsilon)}$, which is more well-behaved and sufficiently close to $\mathcal{H}$.
Constructing the surrogate function class $\mathcal{H}^{(\epsilon)}$.
To construct such a surrogate loss function class $\mathcal{H}^{(\epsilon)}$, for any problem instance $P=(X,y,X_{\text{val}},y_{\text{val}})$, Balcan et al. first approximate the regularized logistic regression estimator $\hat{\beta}{(X,y)}(\lambda)$ by $\beta^{(\epsilon)}{(X,y)}(\lambda)$, calculated using Algorithm 1 if $R(\beta)=\|\beta\|_{1}$ (or Algorithm 2 if $R(\beta)=\|\beta\|_{2}$). Intuitively, for a sufficiently small $\epsilon$, the search space $[\lambda_{\text{min}},\lambda_{\text{max}}]$ can be divided into $\frac{\lambda_{\text{max}}-\lambda_{\text{min}}}{\epsilon}$ pieces. Within each piece, $\beta^{(\epsilon)}{(X,y)}(\lambda)$ is a linear function of $\lambda$, and there exists a uniform error bound of $O(\epsilon^{2})$ for the approximation error $\|\beta^{(\epsilon)}{(X,y)}(\lambda)-\hat{\beta}_{(X,y)}(\lambda)\|_{2}$ (formalized in Theorem D.1). The surrogate validation loss function class can now be defined as
$\mathcal{H}^{(\epsilon)}=\{h^{(\epsilon)}_{\lambda}:\Pi_{m,p}\rightarrow\mathbb{R}_{+}\mid\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]\},$
where
$h^{(\epsilon)}_{\lambda}(P)=\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}\log(1+\exp(-y_{i}x_{i}^{\top}\beta^{(\epsilon)}_{(X,y)}(\lambda)))=\frac{1}{m^{\prime}}\log\left(\prod_{i=1}^{m^{\prime}}(1+\exp(-y_{i}x_{i}^{\top}\beta^{(\epsilon)}_{(X,y)}(\lambda)))\right),$
and in each region $[\lambda_{\text{min}}+t\epsilon,\lambda_{\text{min}}+(t+1)\epsilon]$, $\beta^{(\epsilon)}_{(X,y)}(\lambda))=\lambda a_{t}+b_{t}$, where $a_{t}$, $b_{t}$ is defined in Theorem D.1.
Analyzing $\mathcal{H}^{(\epsilon)}$.
We can now proceed to analyze $\mathcal{H}^{(\epsilon)}$. Moreover, we can simplify the analysis by analyzing the pseudo-dimension of $\mathcal{G}^{(\epsilon)}=\{g^{(\epsilon)}_{\lambda}:\Pi_{m,p}\rightarrow\mathbb{R}_{\geq 0}\mid\lambda_{\min}\lambda_{\max}]\}$, where
$g^{(\epsilon)}_{\lambda}(P)=\prod_{i=1}^{m^{\prime}}(1+\exp(-y_{i}x_{i}^{\top}\beta^{(\epsilon)}_{(X,y)}(\lambda))).$
Fixing a problem instance $P$ and a threshold $\tau\in\mathbb{R}$, by the property of $\beta^{(\epsilon)}_{(X,y)}(\lambda))$, we know that the calculation of ${\text{sign}}(g^{(\epsilon)}_{\lambda}(P)-\tau)$ can be described by $\frac{\lambda_{\max}-\lambda_{\min}}{\epsilon}$ boundary functions and $\frac{\lambda_{\max}-\lambda_{\min}}{\epsilon}$ piece functions. In each interval $[\lambda_{\min}+t\epsilon,\lambda_{\min}+(t+1)\epsilon]$, $g_{\lambda}^{(\epsilon)}(P)$ takes the form
$g^{(\epsilon)}_{\lambda}(P)=\prod_{i=1}^{m^{\prime}}(1+\exp(-y_{i}x_{i}^{\top}(\lambda a_{t}+b_{t}))).$
This implies that, in each interval $[\lambda_{\min}+t\epsilon,\lambda_{\min}+(t+1)\epsilon]$, $g^{(\epsilon)}_{\lambda}(P)$ is a Pfaffian function of degree at most $m$ from a Pfaffian chain $\mathcal{C}$ of length at most $m$ and of Pfaffian degree at most $m$. From Lemma 5.3, we conclude that $\text{Pdim}(\mathcal{H}^{(\epsilon)})=\text{Pdim}(\mathcal{G}^{(\epsilon)})=O(m^{2}+\log(1/\epsilon)).$
Recover the guarantee for $\mathcal{H}$.
Note that we just have the learning guarantee for the approximation function class $\mathcal{H}^{(\epsilon)}$. To recover the learning guarantee for $\mathcal{H}$, we need to leverage the approximation error guarantee between $\mathcal{H}$ and $\mathcal{H}^{(\epsilon)}$ and the following theorem.
Set $\beta^{(\epsilon)}_{0}=\hat{\beta}_{(X,y)}(\lambda_{\min})$, $t=0$, small constant $\delta\in\mathbb{R}_{>0}$, and $\mathcal{A}=\{j\mid[\hat{\beta}_{(X,y)}(\lambda_{\min})]_{j}\neq 0\}.$;
while $\lambda_{t}<\lambda_{\max}$ do
       $\lambda_{t+1}=\lambda_{t}+\epsilon$ ;
      $\left(\beta^{(\epsilon)}_{t+1}\right)_{\mathcal{A}}=\left(\beta_{t}^{(\epsilon)}\right)_{\mathcal{A}}-\left[\nabla^{2}l\left(\beta_{t}^{(\epsilon)},(X,y)\right)_{\mathcal{A}}\right]^{-1}\cdot\left[\nabla l\left(\beta_{t}^{(\epsilon)},(X,y)\right)_{\mathcal{A}}+\lambda_{t+1}\operatorname{sgn}\left(\beta_{t}^{(\epsilon)}\right)_{\mathcal{A}}\right]$;
      $\left(\beta^{(\epsilon)}_{t+1}\right)_{-\mathcal{A}}=\vec{0}$;
      $\mathcal{A}=\mathcal{A}\cup\{j\neq\mathcal{A}\mid\nabla l(\beta_{t+1}^{(\epsilon)},(X,y))>\lambda_{t+1}\}$;
      $\mathcal{A}=\mathcal{A}\setminus\{j\in\mathcal{A}\mid\left|\beta_{t+1,j}^{(\epsilon)}\right|<\delta\}$;
      $t=t+1$;
Algorithm 1 Approximate incremental quadratic algorithm for RLR with $\ell_{1}$ penalty, [Ros04]
Set $\beta^{(\epsilon)}_{0}=\hat{\beta}_{(X,y)}(\lambda_{\min})$, $t=0$.
while $\lambda_{t}<\lambda_{\max}$ do
       $\lambda_{t+1}=\lambda_{t}+\epsilon$ ;
      $\beta^{(\epsilon)}(\lambda)=\beta_{t}^{(\epsilon)}-\left[\nabla^{2}l\left(\beta_{t}^{(\epsilon)},(X,y)\right)+2\lambda_{t+1}I\right]^{-1}\cdot\left[\nabla l\left(\beta_{t}^{(\epsilon)},(X,y)\right)+2\lambda_{t+1}\beta_{t}^{(\epsilon)}\right]$;
      $t=t+1$
Algorithm 2 Approximate incremental quadratic algorithm for RLR with $\ell_{2}$ penalty, [Ros04]
Theorem D.1 (Theorem 1, [Ros04], [BNS24]).
Given a problem instance $P=(X,y,X_{\text{val}},y_{\text{val}})\in\Pi_{m,p}$, for small enough $\epsilon$, if we use Algorithm 1 (2) to approximate the solution $\hat{\beta}_{(X,y)}(\lambda)$ of RLR under $\ell_{1}$ ($\ell_{2}$) constraint by $\beta^{(\epsilon)}_{(X,y)}(\lambda)$ then there is a uniform bound $O(\epsilon^{2})$ on the error $\|\hat{\beta}_{(X,y)}(\lambda)-\beta^{(\epsilon)}_{(X,y)}(\lambda)\|_{2}$ for any $\lambda\in[\lambda_{\min},\lambda_{\max}]$.
For any $\lambda\in[\lambda_{t},\lambda_{t+1}]$, where $\lambda_{k}=\lambda_{\min}+k\epsilon$, the approximate solution $\beta^{(\epsilon)}(\lambda)$ is calculated by
$\beta^{(\epsilon)}_{(X,y)}(\lambda)=\beta_{t}^{(\epsilon)}-\left[\nabla^{2}l\left(\beta_{t}^{(\epsilon)},(X,y)\right)_{\mathcal{A}}\right]^{-1}\cdot\left[\nabla l\left(\beta_{t}^{(\epsilon)},(X,y)\right)_{\mathcal{A}}+\lambda\operatorname{sgn}\left(\beta_{t}^{(\epsilon)}\right)_{\mathcal{A}}\right]=a_{t}\lambda+b_{t},$
if we use Algorithm 1 for RLR under $\ell_{1}$ constraint, or
$\beta^{(\epsilon)}_{(X,y)}(\lambda)=\beta_{t}^{(\epsilon)}-\left[\nabla^{2}l\left(\beta_{t}^{(\epsilon)},(X,y)\right)+2\lambda_{t+1}I\right]^{-1}\cdot\left[\nabla l\left(\beta_{t}^{(\epsilon)},(X,y)\right)+2\lambda\beta_{t}^{(\epsilon)}\right]=a^{\prime}_{t}\lambda+b^{\prime}_{t},$
if we use Algorithm 2 for RLR under $\ell_{2}$ constraint.
Remark 6.
Compared to [BNS24], our framework provides a significant generalization. The techniques used for analyzing the surrogate logistic loss in [BNS24] involves bounding the complexity of Pfaffian piece functions, while the piece boundaries of the surrogate function are nice linear functions. In contrast, in this work we develop general techniques that can handle Pfaffian boundary functions as well, and we recover results from [BNS24] as a special case.
Lemma D.2 ([BSV20]).
Let =$\mathcal{F}=\{f_{r}\mid r\in\mathcal{R}\}$ and $\mathcal{G}=\{g_{r}\mid r\in\mathcal{R}\}$ consist of function mapping $\mathcal{X}$ to $[0,1]$. For any $\mathcal{S}\subseteq\mathcal{X}$, ^$\hat{\mathscr{R}}(\mathcal{F},\mathcal{S})\leq\hat{\mathscr{R}}(\mathcal{G},\mathcal{S})+\frac{1}{\left|\mathcal{S}\right|}\sum_{x\in\mathcal{S}}\|f_{x}^{*}-g_{x}^{*}\|$, where $\hat{\mathscr{R}}$ denotes the empirical Rademacher complexity.
Combining with Lemma A.1, we then recover the following theorem, which establishes the learning guarantee for $\mathcal{H}$.
Theorem D.3 ([BNS24]).
Consider the problem of regularized logistic regression with $\ell_{1}$ (or $\ell_{2}$) regularization. Let $\mathcal{H}$ the validation loss function class, then we have $\mathscr{R}_{n}(\mathcal{H})=O\left(\sqrt{\frac{m^{2}+\log(1/\epsilon)}{n}}+\epsilon^{2}\right)$, where $\mathscr{R}_{n}$ is the Rademacher complexity.
Appendix E Additional background and omitted proofs for Section 7
We record here some fundamental results from the theory of Pfaffian functions (also known as Fewnomial theory) which will be needed in establishing our online learning results. The following result is a Pfaffian analogue of Bezout’s theorem from algebraic geometry, useful in bounding the multiplicity of intersections of Pfaffian hypersurfaces.
Theorem E.1 ([Kho91, Gab95]).
Consider a system of equations $g_{1}({\bm{x}})=\dots=g_{n}({\bm{x}})=0$ where $g_{i}({\bm{x}})=P_{i}({\bm{x}},f_{1}({\bm{x}}),\dots,f_{q}({\bm{x}}))$ is a polynomial of degree at most $d_{i}$ in ${\bm{x}}\in\mathbb{R}^{n}$ and $f_{1},\dots,f_{q}$ are a sequence of functions that constitute a Pfaffian chain of length $q$ and Pfaffian degree at most $M$. Then the number of non-degenerate solutions of this system does not exceed
$2^{q(q-1)/2}d_{1}\dots d_{n}(\min\{q,n\}M+d_{1}+\dots+d_{n}-n+1)^{q}.$
Below, we provide a detailed proof of Theorem 7.2.
See 7.2
Proof.
The proof is similar to analogous results in [BDP20, BS21]. The main difference is that instead of relating the number of ways intervals can label vectors of discontinuity points to the VC-dimension of intervals, we instead relate the number of ways line segments can label vectors of $K$ Pfaffian hypersurfaces of bounded complexity to the VC-dimension of line segments (when labeling Pfaffian hypersurfaces), which from Theorem 7.1 is constant. To verify dispersion, we need a uniform-convergence bound on the number of Lipschitz failures between the worst pair of points $\alpha,\alpha^{\prime}$ at distance $\leq\epsilon$, but the definition allows us to bound the worst rate of discontinuities along any path between $\alpha,\alpha^{\prime}$ of our choice. We can bound the VC dimension of axis aligned segments against bounded complexity Pfaffian hypersurfaces, which will allow us to establish dispersion by considering piecewise axis-aligned paths between points $\alpha$ and $\alpha^{\prime}$.
Let $\mathcal{P}$ denote the set of all Pfaffian hypersurfaces of degree $d$, from a Pfaffian chain of length at most $q$ and Pfaffian degree at most $M$. For simplicity, we assume that every function has its discontinuities specified by a collection of exactly $K$ Pfaffian hypersurfaces ($K$ could be an upper bound, and we could simply duplicate hypersurfaces as needed which does not affect our argument below). For each function $l_{t}$, let $\pi^{(t)}\in\mathcal{P}^{K}$ denote the ordered tuple of Pfaffian hypersurfaces in $\mathcal{P}$ whose entries are the discontinuity locations of $l_{t}$. That is, $l_{t}$ has discontinuities along $(\pi^{(t)}_{1},\dots,\pi^{(t)}_{K})$, but is otherwise $L$-Lispchitz.
For any axis aligned path $s$, define the function $f_{s}:\mathcal{P}^{K}\rightarrow\{0,1\}$ by
[…]
Now, Theorem 7.1 implies that VC dimension of $\mathcal{F}$ is at most $K_{M,q,d,p}$. A standard VC-dimension uniform convergence argument for the class $\mathcal{F}$ imply that with probability at least $1-\delta$, for all $f_{s}\in\mathcal{F}$
$\left\lvert\frac{1}{T}\sum_{t=1}^{T}f_{s}(\pi^{(t)})-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}f_{s}(\pi^{(t)})\right]\right\rvert$ $\leq O\left(\sqrt{\frac{K_{M,q,d,p}+\log(1/\delta)}{T}}\right)\text{, or}$
$\left\lvert\sum_{t=1}^{T}f_{s}(\pi^{(t)})-\mathbb{E}\left[\sum_{t=1}^{T}f_{s}(\pi^{(t)})\right]\right\rvert$ $\leq\tilde{O}\left(\sqrt{T\log(1/\delta)}\right).$
Now since $D(T,s)=\sum_{t=1}^{T}f_{s}(\pi^{(t)})$, we have for all $s$ and $\delta$, with probability at least $1-\delta$, $\sup_{s\in L}D(T,s)\leq\sup_{s\in L}\mathbb{E}[D(T,s)]+\tilde{O}(\sqrt{T\log(1/\delta)})$. Taking expectation and setting $\delta=1/\sqrt{T}$ implies
$\mathbb{E}[\sup_{s\in L}D(T,s)]$ $\leq(1-\delta)\left(\sup_{s\in L}\mathbb{E}[D(T,s)]+\tilde{O}(\sqrt{T\log(1/\delta)})\right)+\delta\cdot T$
$\leq\sup_{s\in L}\mathbb{E}[D(T,s)]+\tilde{O}(\sqrt{T\log(\sqrt{T})})+\sqrt{T},$
which implies the desired bound. Here we have upper bounded the expected regret by $T$ in the low probability failure event where the uniform convergence does not hold. ∎
E.1 Online learning for data-driven agglomerative hierarchical clustering
We will now present some useful lemmas for establishing Theorem 7.3. The following lemma generalizes Lemma 25 of [BDP20].
Lemma E.2.
Let $X_{1},\dots,X_{n}$ be a finite collection of independent random variables each having densities upper bounded by $\kappa$. The random variable $Y=\sum_{i=1}^{n}\beta_{i}X_{i}$ for some fixed scalars $\beta_{1},\dots,\beta_{n}$ with $\sum_{i=1}^{n}\beta_{i}=1$ has density $f_{Y}$ satisfying $f_{Y}(y)\leq\kappa$ for all $y$.
Proof.
We proceed by induction on $n$. First consider $n=1$. Clearly $Y=X_{1}$ and the conclusion follows from the assumption that $X_{1}$ has density upper bounded by $\kappa$.
Now suppose $n>1$. We have $Y=\beta_{1}X_{1}+\beta_{2}X_{2}+\dots+\beta_{n}X_{n}=\beta_{1}X_{1}+(1-\beta_{1})X^{\prime}$, where $X^{\prime}=\frac{1}{1-\beta_{1}}\sum_{i=2}^{n}\beta_{i}X_{i}$. By the inductive hypothesis, $X^{\prime}$ has a density $f_{X^{\prime}}$ which is upper bounded by $\kappa$. Let $f_{X_{1}}$ denote the density of $X_{1}$. Noting $X^{\prime}=\frac{Y-\beta_{1}X_{1}}{1-\beta_{1}}$ and using the independence of $X^{\prime}$ and $X_{1}$, we get
[…]
This completes the induction step. ∎
We will now present a useful algebraic lemma for establishing Theorem 7.3.
Lemma E.3.
If $\alpha>0$, $a,b,c>0$, and $a^{\alpha}+b^{\alpha}-c^{\alpha}>0$, then
$\frac{a^{\alpha}\ln a+b^{\alpha}\ln b-c^{\alpha}\ln c}{a^{\alpha}+b^{\alpha}-c^{\alpha}}\leq\frac{1}{\alpha}+\ln\max\{a,b,c\}.$
Proof.
We consider two cases. First suppose that $c>a$ and $c>b$. We have that
$\frac{a^{\alpha}\ln a+b^{\alpha}\ln b-c^{\alpha}\ln c}{a^{\alpha}+b^{\alpha}-c^{\alpha}}\leq\frac{a^{\alpha}\ln c+b^{\alpha}\ln c-c^{\alpha}\ln c}{a^{\alpha}+b^{\alpha}-c^{\alpha}}\leq\ln c$
in this case.
Now suppose $c\leq a$ (the case $c\leq b$ is symmetric). Observe that for $\alpha>0$, the function $f(x)=x^{\alpha}\ln\frac{K}{x}$ is monotonically increasing for $x\leq Ke^{-1/\alpha}$ when $K,\alpha,x>0$. This implies for $K=\max\{a,b\}e^{1/\alpha}$,
$c^{\alpha}\ln\frac{K}{c}\leq a^{\alpha}\ln\frac{K}{a}\leq a^{\alpha}\ln\frac{K}{a}+b^{\alpha}\ln\frac{K}{b},$
or, equivalently,
$a^{\alpha}\ln a+b^{\alpha}\ln b-c^{\alpha}\ln c\leq\ln K(a^{\alpha}+b^{\alpha}-c^{\alpha}).$
Since, $a^{\alpha}+b^{\alpha}-c^{\alpha}>0$, we further get
$\frac{a^{\alpha}\ln a+b^{\alpha}\ln b-c^{\alpha}\ln c}{a^{\alpha}+b^{\alpha}-c^{\alpha}}\leq\ln K=\frac{1}{\alpha}+\ln\max\{a,b\}.$
Theorem 7.3 (restated).
Consider an adversary choosing a sequence of clustering instances where the $t$-th instance has a symmetric distance matrix $D^{(t)}\in[0,R]^{n\times n}$ and for all $i\leq j,d^{(t)}_{ij}$ is $\kappa$-smooth. For the sub-family of $\mathcal{C}_{1}$ with $\alpha>\alpha_{\min}$ for $0<\alpha_{\min}<\infty$, we have that the corresponding sequence of utility functions $u_{1}^{(t)}$ are $1/2$-dispersed.
Proof.
As noted in the proof of Theorem 6.1, the boundaries of the piecewise-constant dual utility functions are given by
$(d(a_{1},b_{1}))^{\alpha}+(d(a_{2},b_{2}))^{\alpha}=(d(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha}+(d(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha},$ (1)
for some (not necessarily distinct) points $a_{1},b_{1},a_{2},b_{2},a_{1}^{\prime},b_{1}^{\prime},a_{2}^{\prime},b_{2}^{\prime}\in S$.
We use the generalized Descartes’ rule of signs [Jam06, HT11], which implies that the number of real zeros in $\alpha$ (for which boundary condition is satisfied) is at most the number of sign changes when the base of the exponents are arranged in an ascending order (since, the family $\{a^{x}\mid a\in\mathbb{R}_{+}\}$ is Descartes admissible on $\mathbb{R}$), to conclude that the boundary of the loss function occurs for at most one point $\alpha^{*}\in\mathbb{R}$ apart from $\alpha=0$. We consider the following distinct cases (up to symmetry):
  • Case $d_{\beta}(a_{1},b_{1})\geq d_{\beta}(a_{2},b_{2})\geq d_{\beta}(a_{1}^{\prime},b_{1}^{\prime})\geq d_{\beta}(a_{2}^{\prime},b_{2}^{\prime})$. The number of sign changes is one, and the conclusion is immediate.
  • Case $d_{\beta}(a_{1},b_{1})\geq d_{\beta}(a_{1}^{\prime},b_{1}^{\prime})\geq d_{\beta}(a_{2},b_{2})\geq d_{\beta}(a_{2}^{\prime},b_{2}^{\prime})$. The only possibility is $d_{\beta}(a_{1},b_{1})=d_{\beta}(a_{1}^{\prime},b_{1}^{\prime})$ and $d_{\beta}(a_{2},b_{2})=d_{\beta}(a_{2}^{\prime},b_{2}^{\prime})$. But then, the condition holds for all $\alpha$ and we do not get a critical point. This corresponds to the special case of tie-breaking when merging clusters, and we assume ties are broken following some arbitrary but fixed ordering of the pair of points.
  • Case $d_{\beta}(a_{1}^{\prime},b_{1}^{\prime})\geq d_{\beta}(a_{1},b_{1})\geq d_{\beta}(a_{2},b_{2})\geq d_{\beta}(a_{2}^{\prime},b_{2}^{\prime})$. The number of sign changes is two. Since $\alpha=0$ is a solution, there is at most one $\alpha^{*}\in\mathbb{R}\in\setminus\{0\}$ corresponding to a critical point.
Now, let $\epsilon>0$. Consider an interval $I=[\alpha_{1},\alpha_{2}]$ with $\alpha_{1}>\alpha_{\min}$ such that the width $\alpha_{2}-\alpha_{1}\leq\epsilon$. If equation (1) has a solution $\alpha^{*}\in I$ (guaranteed to be unique if it exists), we have that
$d(a_{1},b_{1})=\left((d(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha^{*}}+(d(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha^{*}}-(d(a_{2},b_{2}))^{\alpha^{*}}\right)^{1/{\alpha^{*}}}$
Let $p_{1}$ denote the density of $d(a_{1},b_{1})$ which is at most $\kappa$ by assumption, and let $f(\alpha^{*}):=(d(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha^{*}}+(d(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha^{*}}-(d(a_{2},b_{2}))^{\alpha^{*}}$ and $d_{1}(\alpha^{*}):=\left(f(\alpha^{*})\right)^{1/{\alpha^{*}}}$. We seek to upper bound $\Pr_{p_{1}}\{d_{1}(\alpha^{*})\mid\alpha^{*}\in I,f(\alpha^{*})\geq 0\}$ to get a bound on the probability that there is a critical point in $I$.
Now we consider two cases. If $f(\alpha)\leq\epsilon^{{\alpha}}$ for all $\alpha\in I$, we have that $d_{1}(\alpha)=f(\alpha)^{1/\alpha}\leq\epsilon$ for all $\alpha\in I$. Thus, the probability of having a critical point in $I$ is $\mathcal{O}(\kappa\epsilon)$ in this case.
Else, we have that $f(\alpha^{*})>\epsilon^{\alpha^{*}}>0$ for some $\alpha^{*}\in I$. Using Taylor series expansion for $d_{1}(\alpha)$ around $\alpha^{*}$, we have
$d_{1}(\alpha^{*}+\epsilon)=d_{1}(\alpha^{*})+d_{1}^{\prime}(\alpha^{*})\epsilon+\mathcal{O}(\epsilon^{2}).$
Thus, the set $\{d_{1}(\alpha^{*})\mid\alpha^{*}\in I,f(\alpha^{*})\geq 0\}$ is contained in an interval of width at most $|d_{1}^{\prime}(\alpha^{*})|\mathcal{O}(\epsilon)$, implying a upper bound of $\kappa|d_{1}^{\prime}(\alpha^{*})|\mathcal{O}(\epsilon)$ on the probability of having a critical point in $I$.
Thus it is sufficient to give a bound on $|d_{1}^{\prime}(\alpha^{*})|$. We have
$d_{1}^{\prime}(\alpha^{*})$ $=d_{1}(\alpha^{*})\left(-\frac{1}{{\alpha^{*}}^{2}}\ln f(\alpha^{*})+\frac{1}{\alpha^{*}}\frac{g(\alpha^{*})}{f(\alpha^{*})}\right)$
$=\frac{1}{\alpha^{*}}\left(\frac{g(\alpha^{*})}{f(\alpha^{*})}-d_{1}(\alpha^{*})\ln d_{1}(\alpha^{*})\right),$
where
$g(\alpha):=\ln d(a_{1}^{\prime},b_{1}^{\prime})(d(a_{1}^{\prime},b_{1}^{\prime}))^{\alpha^{*}}+\ln d(a_{2}^{\prime},b_{2}^{\prime})(d(a_{2}^{\prime},b_{2}^{\prime}))^{\alpha^{*}}-\ln d(a_{2},b_{2})(d(a_{2},b_{2}))^{\alpha^{*}}.$
Thus,
$|d_{1}^{\prime}(\alpha^{*})|$ $=\Bigg{\lvert}\frac{1}{\alpha^{*}}\left(\frac{g(\alpha^{*})}{f(\alpha^{*})}-d_{1}(\alpha^{*})\ln d_{1}(\alpha^{*})\right)\Bigg{\rvert}$
$\leq\Big{\lvert}\frac{1}{\alpha^{*}}\Big{\rvert}\cdot\left(\Big{\lvert}\frac{g(\alpha^{*})}{f(\alpha^{*})}\Big{\rvert}+\lvert d_{1}(\alpha^{*})\ln d_{1}(\alpha^{*})\rvert\right)$
$\leq\Big{\lvert}\frac{1}{\alpha_{\min}}\Big{\rvert}\cdot\left(\Big{\lvert}\frac{g(\alpha^{*})}{f(\alpha^{*})}\Big{\rvert}+R\ln R\right).$
Now using Lemma E.3, we get that
$\Big{\lvert}\frac{g(\alpha^{*})}{f(\alpha^{*})}\Big{\rvert}\leq\frac{1}{\alpha^{*}}+\ln R,$
and thus,
$|d_{1}^{\prime}(\alpha^{*})|\leq O\left(\frac{R\ln R}{\alpha_{\min}^{2}}\right).$
Using that we have at most $n^{8}$ boundary functions of the form (1), we can conclude that $\mathbb{E}[D(T,s)]=\tilde{O}(\frac{R\ln R}{\alpha_{\min}^{2}}\kappa n^{8}\epsilon T)$ and using Theorem 7.2 we can conclude that the sequence of utility functions are $\frac{1}{2}$-dispersed.
To establish Theorem 7.4, we first present a useful lemma, and restate a useful result from [BS21].
Lemma E.4.
Suppose $X$ is a real-valued random variable taking values in $[0,M]$ for some $M\in\mathbf{R}^{+}$ and suppose its probability density is $\kappa$-bounded. Then, $Y=X^{\alpha}$ for $\alpha\in[\alpha_{\min},1]$ for some $\alpha_{\min}>0$ takes values in $[0,M]$ with a $\kappa^{\prime}$-bounded density with $\kappa^{\prime}\leq\frac{\kappa}{\alpha_{\min}}\max\{1,M^{\frac{1}{\alpha_{\min}}-1}\}$.
Proof.
The cumulative density function for $Y$ is given by
$F_{Y}(y)$ $=\Pr[Y\leq y]=\Pr[X\leq y^{1/\alpha}]$
[…]
where $f_{X}(x)$ is the probability density function for $X$. Using Leibniz’s rule, we can obtain the density function for $Y$ as
$f_{Y}(y)$ $=\frac{d}{dy}F_{Y}(y)$
[…]
$\leq\kappa\frac{d}{dy}y^{1/\alpha}$
$=\frac{\kappa}{\alpha}y^{\frac{1}{\alpha}-1}.$
Now for $y\leq 1$, $y^{\frac{1}{\alpha}-1}\leq y\leq 1$ and therefore $f_{Y}(y)\leq\frac{\kappa}{\alpha_{\min}}$. Else, $y\leq M$, and $f_{Y}(y)\leq\frac{\kappa}{\alpha_{\min}}M^{\frac{1}{\alpha_{\min}}-1}$. ∎
We will also need the following result which is useful to establish dispersion when the discontinuities of the loss function are given by roots of an exponential equation in the parameter with random coefficients.
Theorem E.5 ([BS21]).
Let $\phi(x)=\sum^{n}_{i=1}a_{i}e^{b_{i}x}$ be a random function, such that coefficients $a_{i}$ are real and of magnitude at most $R$, and distributed with joint density at most $\kappa$. Then for any interval $I$ of width at most $\epsilon$, $\Pr(\phi\text{ has a zero in }I)\leq\tilde{O}(\epsilon)$ (dependence on $b_{i},n,\kappa,R$ suppressed).
We will now re-state and prove our online learning result for learning the parameter $\alpha$ in the clustering linkage family $\mathcal{C}_{3}$.
Theorem 7.4 (restated).
Consider an adversary choosing a sequence of clustering instances where the $t$-th instance has a symmetric distance matrix $D^{(t)}\in[0,R]^{n\times n}$ and for all $i\leq j,d^{(t)}_{ij}$ is $\kappa$-smooth. For the family of clustering algorithms $\mathcal{C}_{3}$, we have that the corresponding sequence of utility functions $u_{3}^{(t)}$ as a function of the parameter $\alpha=(\alpha_{1},\dots,\alpha_{L})$ are $1/2$-dispersed.
Proof.
We will show dispersion for each $\alpha_{i}$ keeping the remaining parameters fixed. This is sufficient because, the definition of dispersion [BDP20] allows us to consider discontinuities along axis-aligned paths between pairs of points $\alpha,\alpha^{\prime}$ in the parameter space. WLOG, assume that $\alpha_{2},\dots,\alpha_{L}$ are fixed. The boundary functions are given by solutions of exponential equations in $\alpha_{1}$ of the form
$\frac{1}{|A||B|}\sum_{a\in A,b\in B}\Pi_{i\in[L]}(d_{i}(a,b))^{\alpha_{i}}=\frac{1}{|A^{\prime}||B^{\prime}|}\sum_{a^{\prime}\in A^{\prime},b^{\prime}\in B^{\prime}}\Pi_{i\in[L]}(d_{i}(a^{\prime},b^{\prime}))^{\alpha_{i}},$
for $A,B,A^{\prime},B^{\prime}\subseteq S$. We can rewrite this equation in the form $\sum^{n}_{j=1}a_{j}e^{b_{j}x}=0$ where, $x=\alpha_{1}$, $b_{j}=\ln(d_{1}(a,b))$ for $a,b\in A\times B$ or $a,b\in A^{\prime}\times B^{\prime}$ and $a_{j}=\frac{1}{|A||B|}\Pi_{i\in\{2,\dots,L}d_{i}(a,b)^{\alpha_{i}}$ for $a,b\in A\times B$ or $a_{j}=\frac{-1}{|A^{\prime}||B^{\prime}|}\Pi_{i\in\{2,\dots,L\}}d_{i}(a^{\prime},b^{\prime})^{\alpha_{i}}$ for $a^{\prime},b^{\prime}\in A^{\prime}\times B^{\prime}$. The coefficients $a_{j}$ are real with magnitude at most $R^{L-1}$. By Lemma E.4, we have that $d_{i}(a,b)^{\alpha_{i}}$ is $\kappa^{\prime}$-bounded for $\kappa^{\prime}\leq\frac{\kappa}{\alpha_{\min}}\max\{1,R^{\frac{1}{\alpha_{\min}}-1}\}$, and therefore the coefficients $a_{j}$ have a $\kappa^{\prime\prime}$-bounded density for $\kappa^{\prime\prime}\leq n^{2}\kappa^{\prime L-1}=O\left(n^{2}\left(\frac{\kappa R^{1/\alpha_{\min}}}{\alpha_{\min}}\right)^{L-1}\right)$ (using Lemma 8 from [BDV18]). Using Theorem E.5, the probability there is a discontinuity along any segment along the direction $\alpha_{1}$ of width $\epsilon$ is $p_{1}=\tilde{O}(\epsilon)$. Thus, for any axis-aligned path $s$ between points $\alpha,\alpha^{\prime}\in\mathbf{R}^{L}$, the expected number of discontinuities for any utility function $u_{3}^{(t)}$ ($t\in[T]$) is at most $Lp_{1}=\tilde{O}(\epsilon)$. We can now apply Theorem 7.2 to get
$\mathbb{E}[\sup_{s\in\mathcal{L}}D(T,s)]$ $\leq\sup_{s\in\mathcal{L}}\mathbb{E}[D(T,s)]+\tilde{O}(\sqrt{T\log T})$
$=\tilde{O}(\epsilon T)+\tilde{O}(\sqrt{T\log T}),$
where $\mathcal{L}$ denotes the set of axis-aligned paths between pairs of points $\alpha,\alpha^{\prime}\in\mathbf{R}^{L}$. It then follows that $\mathbb{E}[\sup_{s\in\mathcal{L}}D(T,s)]=\tilde{O}(\sqrt{T})$ for $\epsilon\geq T^{-1/2}$, establishing that the sequence of utility functions $u_{3}^{(1)},\dots,u_{3}^{(T)}$ is $1/2$-dispersed. ∎
E.2 Online learning for data-driven regularized logistic regression
Theorem 7.5 (restated).
Consider the online learning problem for tuning the logistic regression regularization coefficient $\lambda_{t}$ stated above. There exists an online learner with expected regret bounded by $\mathcal{O}(\sqrt{T\log[(\lambda_{\text{max}}-\lambda_{\text{min}})T]})$.
Proof.
We consider an $\epsilon$-grid of $\lambda$ values given by intervals $[\lambda_{\text{min}}+k\epsilon,\lambda_{\text{min}}+(k+1)\epsilon]$ for $k=0,\dots,\lfloor\frac{\lambda_{\max}-\lambda_{\min}}{\epsilon}\rfloor$. For each round $t$, to construct the surrogate loss function, we sample a uniformly random point $\lambda_{t}^{(k)}$ from each interval $[\lambda_{\text{min}}+k\epsilon,\lambda_{\text{min}}+(k+1)\epsilon]$ and compute the surrogate model $\hat{\beta}_{(X,y)}(\lambda)$ at that point using Algorithm 1 (2) for $\ell_{1}$ ($\ell_{2}$), and use a linear interpolation between successive points $\lambda_{t}^{(k)},\lambda_{t}^{(k+1)}$ which are at most $2\epsilon$ apart. By Theorem D.1, we have that $\|\hat{\beta}_{(X_{t},y_{t})}(\lambda)-\beta^{(\epsilon)}_{(X_{t},y_{t})}(\lambda)\|_{2}=\mathcal{O}(\epsilon^{2})$ for any $\lambda\in[\lambda_{\min},\lambda_{\max}]$, where $\beta^{(\epsilon)}_{(X_{t},y_{t})}(\lambda)$ is the true model (that exactly optimizes the logistic loss) for $(X_{t},y_{t})$. By Lemma 4.2 of [BNS24] we can conclude that the corresponding surrogate loss $\hat{h}^{(\epsilon)}_{\lambda}(P_{t})$ also satisfies $||\hat{h}^{(\epsilon)}_{\lambda}(P_{t})-h_{\lambda}(P_{t})||\leq\mathcal{O}(\epsilon^{2})$ for any $\lambda\in[\lambda_{\min},\lambda_{\max}]$.
This implies the regret with respect to the true loss is at most $\mathcal{O}(\epsilon^{2}T)$ more than the regret when using the surrogate function. Suppose $\lambda^{*}=\operatorname*{arg\,min}_{\lambda\in[\lambda_{\min},\lambda_{\max}]}\sum_{t=1}^{T}\mathbb{E}[h_{\lambda}(P_{t})]$. We have,
$R_{T}$ $=\sum_{t=1}^{T}\mathbb{E}[h_{\lambda_{t}}(P_{t})-h_{\lambda^{*}}(P_{t})]$
$=\sum_{t=1}^{T}\mathbb{E}[h_{\lambda_{t}}(P_{t})-\hat{h}_{\lambda^{*}}(P_{t})]+\sum_{t=1}^{T}\mathbb{E}[\hat{h}_{\lambda^{*}}(P_{t})-{h}_{\lambda^{*}}(P_{t})]$
$\leq\sum_{t=1}^{T}\mathbb{E}[h_{\lambda_{t}}(P_{t})-\hat{h}_{\lambda^{*}}(P_{t})]+\mathcal{O}(\epsilon^{2}T)$
$\leq\sum_{t=1}^{T}\mathbb{E}[h_{\lambda_{t}}(P_{t})-\hat{h}_{\hat{\lambda}}(P_{t})]+\mathcal{O}(\epsilon^{2}T),$
where $\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in[\lambda_{\min},\lambda_{\max}]}\sum_{t=1}^{T}\mathbb{E}[\hat{h}_{\lambda}(P_{t})]$.
We next show that the surrogate function is $f$-point-dispersed (Definition 4, [BDP20]) with $f(T,r)=\frac{rT}{\epsilon}$. Indeed, let $0<r<\epsilon$. Let $I=[\lambda_{1},\lambda_{2}]\subset[\lambda_{\min},\lambda_{\max}]$ be an interval of length $r$. Then the probability $\hat{h}_{\lambda}(P_{t})$ has a critical point in $I$ is at most $r/\epsilon$.
Using Theorem 5 of [BDP20] combined with the above argument, we now get that there is an online learner with regret $\mathcal{O}(\sqrt{T\log((\lambda_{\text{max}}-\lambda_{\text{min}})/r)}+f(T,r)+Tr+\epsilon^{2}T)=\mathcal{O}(\sqrt{T\log((\lambda_{\text{max}}-\lambda_{\text{min}})/r)}+\frac{Tr}{\epsilon}+\epsilon^{2}T)$. Setting $\epsilon=T^{-1/4}$ and $r=T^{-3/4}$, we get the claimed regret upper bound. ∎